### v1

# Mamba State-Space Models Can Be Strong Downstream Learners

Abstract
Mamba [22] state-space models (SSMs) have recently outperformed state-of-the-art (SOTA) Transformer large language models (LLMs) in various tasks and been widely adapted. However, Mamba’s downstream learning capabilities remain either unexplored–e.g., mixed-precision (MPFT) and parameter-efficient fine-tuning (PEFT)–or under-evaluated–e.g., in-context learning (ICL). For the latter, recent works [42, 19] reported Mamba’s ICL rivals SOTA Transformer LLMs using non-standard benchmarks. In contrast, we show that on standard benchmarks, pretrained Mamba models achieve only 38% of the ICL performance improvements (over zero-shot) of comparable Transformers.
Enabling MPFT and PEFT in Mamba architectures is challenging due to recurrent dynamics and highly customized CUDA kernels, respectively. However, we prove that Mamba’s recurrent dynamics are robust to small input changes using dynamical systems theory. Empirically, we show that performance changes in Mamba’s inference and fine-tuning due to mixed-precision align with Transformer LLMs. Furthermore, we show that targeting key memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus achieving parameter efficiency while retaining speedups. We show that combining MPFT and PEFT enables up to 2.15 times more tokens-per-second and 65.5% reduced per-token-memory compared to full Mamba fine-tuning, while achieving up to 81.5% of the ICL performance improvements (over zero-shot) of comparably fine-tuned Transformers.
**footnotetext: Equal Contribution. Correspondence to john.t.halloran@leidos.com.
1 Introduction
Innovating on previous state-space models (SSMs) [23, 11], Mamba [22] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs). Mamba was initially shown to greatly outperform comparable Transformer LLMs [5] across a large number of standard natural language benchmarks. Subsequently, pretrained Mamba models have been widely adapted across different data modalities [39, 62, 34, 43, 35], tasks [57, 59, 45, 60, 54, 35, 2], and architectures [1, 42, 37].
However, despite such rapid and widespread adaptation, evaluation of Mamba’s ability to perform standard downstream learning abilities exhibited by Transformer-based LLMs have either not been extensively conducted on standard natural benchmarks or are completely lacking. For instance, while recent works [42, 19, 29] have evaluated Mamba’s ability to perform in-context learning (ICL), such studies focused extensively on either non-natural tasks [29, 17] or non-standard benchmarks [25].
Furthermore, evaluation of Mamba’s mixed-precision fine-tuning (MPFT) and performance efficient fine-tuning (PEFT) capabilities are currently lacking. For the former, MPFT (and, by extension, mixed-precision inference) are made difficult due to potential sensitivities of Mamba’s recurrent dynamics, where  [21, 28] suggest full precision (FP32) is required to perform stable training. For the latter, PEFT via standard low-rank adaptation (LoRA) [27] is made difficult within Mamba’s SSM layer (referred to herein as the MambaBlock) due to highly customized SSM CUDA kernels which provide competitive performance to attention-based speedups [10] at the cost of standard adapter support. However, PEFT and MPFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [52], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12].
Herein, we extensively explore Mamba’s downstream learning capabilities across standard natural benchmarks. For ICL, we show that, in contrast to recent non-standard studies showing Mamba models rival state-of-the-art (SOTA) LLMs of similar parameter counts, the pretrained benefits of Mamba few-shot learning are significantly less than comparable Transformer LLMs across standard natural benchmarks; averaged across the benchmarks and parameter counts in Table 1, Mamba models only achieve 38% of the performance improvements (relative to zero-shot) of comparable Transformer models from the Pythia suite [5]. However, we show in the sequel that Mamba models can more than halve this gap through efficient fine-tuning, achieving as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers.
For MPFT, we leverage theory from dynamical systems to show that small input changes in a MambaBlock do not lead to exponentially deviating outputs. Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference and fine-tuning are on par with those demonstrated by Transformer LLMs (Section 7). For PEFT, we show that by targeting the largest memory buffer exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying. We show that this leads to extremely efficient PEFT, resulting in up to 2.15 times faster training and 65.5% reduced memory compared to the largest evaluated Mamba model without MPFT or PEFT.
Table 1: In-context learning performance for pretrained Mamba and Pythia models. Models are collected into parameter classes for head-to-head comparison using the groupings in [22]. Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16]. LAMBADA zero-shot is more effective for the model sizes considered (further discussed in [58, 8]) and thus excluded from few-shot performance averages. Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
Model $N$-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
Mamba 130M 0 16.07 44.3 35.3 64.5 48.0 24.2 44.8
1 19.34 38.3 35.2 64.3 47.1 23.5 51.3 -1.4
3 23.13 35.4 35.1 65.1 49.0 24.0 50.7 -0.2
5 24.38 36.2 34.8 64.9 49.2 23.9 50.5 -0.5
[…]
1 47.21 28.2 30.6 62.2 43.4 23.7 49.3 -0.4
3 63.70 24.7 30.5 61.9 44.8 22.9 51.3 0.1
5 66.30 25.3 30.4 62.6 43.4 23.1 50.8 -0.2
Mamba 370M 0 8.14 55.6 46.5 69.5 55.0 27.9 55.5
1 9.74 49.8 45.9 69.3 57.4 26.5 54.6 -0.8
3 10.89 48.5 46.2 69.6 58.7 28.5 53.6 1.0
5 11.36 48.5 46.2 69.4 58.3 28.0 56.0 1.3
[…]
1 12.26 47.1 40.5 68.0 53.8 25.6 52.4 1.8
3 14.39 43.2 40.9 67.9 55.1 26.9 54.0 4.2
5 14.62 44.1 40.8 68.1 54.6 26.6 53.4 3.5
Mamba 790M 0 6.01 61.7 55.1 72.1 61.2 29.6 56.0
1 7.06 56.2 54.5 72.5 63.3 30.1 56.9 1.4
3 8.05 54.8 54.2 72.2 63.4 31.6 57.2 2.4
5 8.83 53.4 54.6 72.5 64.6 32.1 57.5 3.4
[…]
1 8.99 51.8 47.3 70.7 57.1 28.2 53.4 1.0
3 10.48 48.2 47.5 71.2 59.2 28.0 54.3 2.2
5 10.86 48.4 47.3 71.4 58.7 28.4 53.1 1.9
Mamba 1.4B 0 5.04 65.0 59.1 74.2 65.5 32.9 58.6
1 5.83 60.6 58.20 74.7 64.5 33.0 61.2 -0.5
3 6.62 58.9 58.8 73.7 66.1 34.4 60.9 0.6
5 6.98 58.4 59.0 74.0 66.4 35.5 60.5 1.4
[…]
1 6.96 56.3 52.1 71.4 62.0 29.5 57.5 1.4
3 7.89 54.4 52.6 70.9 63.9 31.1 56.8 2.9
5 8.02 54.4 52.8 71.0 63.2 31.3 57.8 3.3
Mamba 2.8B 0 4.23 69.2 66.2 75.2 69.7 36.3 63.4
1 5.01 63.9 65.7 75.5 69.8 37.2 63.7 0.6
3 5.53 63.0 65.5 75.2 70.8 38.1 64.8 1.6
5 5.70 62.7 66.2 76.2 70.9 38.3 64.6 2.1
[…]
1 5.66 60.9 59.4 73.8 66.8 34.8 59.0 1.7
3 6.20 59.1 59.9 74.7 67.4 34.9 60.8 2.9
5 6.52 59.1 60.2 74.5 67.1 35.0 61.3 3.1
2 Background
2.1 Downstream learning for LLMs
Since the release of the Transformer architecture [51], attention-based LLMs have exhibited several downstream learning abilities–in particular, PEFT, MPFT, and ICL–which allow the rapid adaptation of foundation models towards specific applications. PEFT using adapters [24] allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters. Arguably the most widely used such PEFT method is LoRA [27], which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
To further decrease the computational demands necessary for LLM fine-tuning and inference, MPFT via mixed-precision (i.e., FP16 or BF16[30, 40] and quantized low-precision [12] have proven effective strategies to reduce GPU memory and runtime requirements without deleterious effects on downstream performance [12, 56]. Additionally, mixed-precision approaches have paved the way for hardware-aware optimizations within the self-attention module [10], greatly mitigating the quadratic complexity of Transformer LLMs. Together, PEFT and MPFT have created a rich ecosystem with which varying combinations of these approaches may be used to meet the computational constraints of a given training system. We note that post-fine-tuning quantization approaches [13] may be further used to decrease Transformer LLM computational demands, but such approaches are not considered in this work.
ICL provides an adaptable alternative to fine-tuning. Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest. Given sufficiently large models and pretraining data [8, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting. ICL thus avoids computational and storage costs associated with fine-tuning and PEFT adapter weights, respectively, and rather changes the prompt to immediately change the model for downstream learning. However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [8, 38].
2.2 State-space Models
Structured state-space sequence (S4) models [23, 14] are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency. Within the S4 layer, an input signal is discretized and LTI parameters representing the input’s latent dynamics are learned. Owing to the S4 block’s latent dynamics being LTI, the S4 block’s output may be thus compactly represented as a single convolution between the input and an SSM convolution kernel (a matrix whose entries are products of LTI learnable parameters resulting from unrolling the state-space equations). However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures [22].
Innovating on these previous S4 approaches, Mamba utilizes time-varying parameters to model latent dynamics, thus broadening the ability to capture nuanced changes evolving in discrete-time. Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations [14]. To enable hardware efficiency with time-varying SSM parameters, [22] thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states.
3 Mamba state-space models
For model dimension $d$ and maximum input sequence length $T$, the MambaBlock defines state-space parameters $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}$. The matrix $\bm{\Delta}_{t}$ controls the discrete step-size. Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{T}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{T}\in\mathbb{R}{}^{d}$:
$\bm{x}_{t}$ $=\bar{\mathbf{A}}_{t}\bm{x}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{u}_{t}$ (1)
$\mathbf{y}_{t}$ $=\bar{\mathbf{C}}_{t}\bm{x}_{t},$ (2)
where $\bar{\bm{\Delta}}_{t}=\texttt{softplus}(\texttt{Linear}(\bm{\Delta}_{t}))\in\mathbb{R}{}^{d\times d}$, $\bar{\mathbf{A}}_{t}=\exp{(\bar{\bm{\Delta}}_{t}\mathbf{A})}$ and $\bar{\mathbf{B}}_{t}=\mathbf{A}^{-1}(\bar{\mathbf{A}}-\mathbf{I})\mathbf{B}_{t}$. In practice, $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are diagonal matrices.
3.1 Hardware-aware optimizations
As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [11] are no longer applicable. However, by diagonality, each dimension may be computed in parallel. Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [7]. [15] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated. Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{n\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$ (3)
where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on.
4 Stable dynamics in the MambaBlock
The Mamba foundation models were pretrained in full FP32 precision. Consequently, official Mamba implementations have cautioned against fine-tuning or training in reduced precision [21, 28], with potential sensitivities of MambaBlock recurrent dynamics remaining an open question. We answer the latter using theory from dynamical systems. For Mamba’s discrete dynamic system in Equations 1 and  2, define
$\bm{x}_{t}$ $=F_{\theta}(\bm{x}_{t-1},\mathbf{u}_{t}),$ (4)
where $\theta$ denotes the time-varying parameters described in Section 3. For input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ and initial latent state vector $\bm{x}_{0}$, we thus write
$\bm{x}_{T}$ $=F_{\theta}(F_{\theta}(\dots F_{\theta}(\bm{x}_{0},\mathbf{u}_{1})))\coloneq F_{\theta}^{T-1}(\bm{x}_{0},\mathbf{u}_{1}).$
The rate of divergence between two scalar $\varepsilon$-close inputs to a discrete dynamical system is bounded by the system’s maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ [41]. Given $\lambda_{\mathtt{max}}{}$ and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$, the maximum deviation between these points grows as [32, 47]:
$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\lambda_{\mathtt{max}}{})}).$
Thus, when $\lambda_{\texttt{max}}>0$, nearby trajectories exponentially separate and, when $\lambda_{\texttt{max}}\leq 0$, nearby trajectories ultimately converge to the same fixed point or periodic cycles.
The maximal Lyapunov exponent is defined as
$\lambda_{\mathtt{max}}{}$ $\coloneq\lim_{T\to\infty}\frac{1}{T}\log\left\lVert\prod_{t=0}^{T}\frac{\partial\bm{x}_{t}}{\partial\bm{x}_{t-1}}\right\rVert_{2},$
where $\left\lVert\right\rVert_{2}$ denotes the spectral norm for matrices. For an arbitrary MambaBlock, we prove the following:
Theorem 1.
Let $(\bm{x}_{t-1},\mathbf{u}_{t})$ be the latent state and input at an arbitrary time $t\in\{1,\dots,T\}$ within a MambaBlock. Then small changes $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
The proof of Theorem 1 is available in Appendix A, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
4.1 Consequences for automatic mixed-precision
During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16). During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating. In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ fed into the SSM scan (which are passed through a SwiGLU), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
For a discrete dynamical system with $\lambda_{\texttt{max}}>0$, changes due to AMP compound after repeated expansion of the recurrent state, thus leading to exponential deviations between quantities calculated using mixed- versus full-precision. We note that Transformers are not recurrent, and thus not susceptible to such issues. Yet, just as differences introduced by quantization/mixed-precision produce output differences in Transformer results, differences are expected in Mamba results using different precision strategies. However, by Theorem 1, such differences do not exponentially compound over discrete-time within the MambaBlock.
5 Regularized fine-tuning using LoRA in the MambaBlock
Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer. E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these invidiual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer). Rather than adjusting Mamba’s low-level CUDA kernels themselves to integrate LoRA within the highly optimized prefix scan, we note that $\mathbf{W}$ can instead be directly targeted. Doing so, we have the following, where the proof is available in Appendix B.
Theorem 2.
Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 3. Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
Thus, the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
6 Related Work
Several recent works [42, 19, 29, 37] have studied Mamba’s ability to perform ICL. However, none of these have extensively studied Mamba’s ICL capabilities either on standard NLP benchmarks or on pure MambaBlock foundation models. In particular, foundational Mamba models’ ICL abilities were tested in [42] to learn simple function classes (e.g., logistic regression and decision trees [17]) and in  [19] to learn non-standard NLP benchmarks (i.e., task vectors [25]). While [42, 19] report Mamba’s ICL abilities rival SOTA Transformers, their utilized benchmarks were proposed as supplemental ICL studies after Transformer LLMs’ success on standard NLP benchmarks [8]. Indeed, direct evaluation of Mamba foundation models on standard NLP benchmarks does not lead to higher gains over zero-shot performance relative to comparable Transformer LLMs (demonstrated in Table 1).
Lyapunov exponents have previously been considered for classic RNN structures (e.g., vanilla RNNs, LSTMs, GRUs, PLRNNs, etc.) [41, 53], to determine when such models exhibit chaotic dynamics and the impact on the exploding/vanishing gradient phenomena***We note that this continues a long line of research exploring RNNs sensitivity to initial conditions and their subsequent ability to produce chaotic output [44, 33, 3, 4], although previous work did not leverage Lyapunov exponents.. For more recent S4 neural models, [18] used Hurwitz matrices to characterize the numerical stability of linear time-invariant (LTI) S4 models. However, such analysis is not applicable to time-varying models, such as Mamba, nor does it characterize the effects of sensitive dependence on initial conditions (e.g., divergence of two $\varepsilon$ close inputs). To the best of our knowledge, no previous works have used Lyapunov exponents to explore the effects of mixed-precision on recurrent neural models or Mamba architectures.
As in  [22], the majority of subsequent Mamba works have focused on pretraining MambaBlocks using full precision [62, 59, 1, 37]. Notably, the official implementation of Jamba [37], the Transformer-Mamba hybrid, supports mixed- and 8-bit precision, but avoids MambaBlocks when applying such quantization [31]. Similarly, the official Mamba sources advise using full precision within the MambaBlock [28, 21], cautioning against using mixed-precision due to potential recurrent sensitivities. To the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or explored pure Mamba networks fine-tuning abilities relative to Transformer LLMs.
7 Experiments
To demonstrate the implications of Theorem 1, we explore the performance difference between running inference with full-precision pretrained weights and using mixed-precision (FP16 and BF16) weights. Model performance is measured as percent accuracy using the MMLU [26] dataset. The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy. Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy.
Table 2: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference. Divergence is averaged over {0, 1, 3, 5}-shot performance. Pretrained checkpoints are used for Mamba (M), Pythia (P), OLMo [20], and Phi-1.5 [36] (Phi) models.
Model M P M P M P OLMo M P Phi M P
Size 130m 160m 370m 410m  790m 1b 1.4b 1.5b 2.8b
FP16 $\mu$ 0.03 0.35 0.05 0.06 0.21 0.05 0.04 0.04 0.07 0.03 0.15 0.12
BF16 $\mu$ 0.05 1.45 0.20 0.20 0.66 0.16 0.13 0.31 0.13 1.05 1.17 0.11
Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [5], OLMo [20] 336B-token checkpoint, and Phi 1.5 [36]). We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers. To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory). All models were evaluated using the LM evaluation harness from Eleuther AI [16]. Further experimental details are available in Appendix C. The results are available in Table 2.
From Table 2, inferencing in Pythia using FP16 and BF16 result in an average 0.13 and 0.41 full-precision divergence, respectively. Mamba displays similar averages in comparison: inferencing in Mamba using FP16 and BF16 result in an average 0.10 and 0.48 divergence, respectively. Interestingly, both SSM and Transformer architectures exhibit large divergence spikes–i.e., mean divergence greater than a percentage point–when using BF16, which occurs once for Mamba and Phi 1.5 models and twice for Pythia models. In the following, we show that such spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
7.1 Non-divergent Mamba fine-tuning
We next explore the implications of Theorem 1 on fine-tuning, wherein mixed-precision is especially critical; MPFT combined with PEFT adapters have been shown to drastically reduce Transformer fine-tuning times [12]. We are thus interested in the divergence between Mamba models fully fine-tuned (i.e., no adapters, all model weights are trained) in full-precision and models fine-tuned using mixed-precision and/or PEFT adapters. We focus on utilizing LoRA [27], which is arguably the most widely used PEFT framework for LLMs.
Using the Alpaca dataset [48], Mamba 160M, 410M, and 790M models are fine-tuned for three epochs with a maximum sequence length of 512. We denote the targeting of all linear layers (ALL) for LoRA as ALL LoRA, the targeting of a subset of linear layers (SLL) for LoRA as SLL LoRA, and no adapters as Full (i.e., full fine-tuning). Both ALL and SLL LoRA adapt the large memory buffer described in Theorem 2.
Each fine-tuning run occurred on a single A10G GPU. To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size. While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules. For comparison, Pythia 160M, 410M, and 1B models are fine-tuned using the same experimental setup. The training recipe for all models was adapted from [50], with the AdamW_torch optimizer and a cosine annealing schedule. Further experimental details are available in Appendix C.
Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba and Pythia models. Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
For each Mamba and Pythia model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy. Across mixed-precisions and adapter settings, Mamba displays comparable divergences to Pythia models. E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia. Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia. Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning (in contrast to Pythia models).
7.2 Hardware throughput and memory-utilization improvements given PEFT and MPFT
With comparable divergences to Transformers and stable dynamics, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs. To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset. However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
For each MPFT and PEFT configuration, the average tokens-per-second (ATPS) is calculated as the total tokens used for fine-tuning divided by total training time, and the maximum memory-per-token (MMPT) is calculated as the maximum GPU memory utilization incurred (over the entire fine-tuning run) divided by the total number of tokens in each mini-batch. Results are plotted in Figure 2(b).
(a) Average tokens-per-second $\uparrow$.
(b) Maximum memory-per-token $\downarrow$.
Figure 2: Timing and memory usage calculated Mamba model-sizes and PEFT combinations. Each model was trained using the Alpaca dataset [48] dataset for three epochs and maximum sequence length 512. For each PEFT combination, the batch size was tuned to maximize GPU occupancy.
Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 2(b). Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%. Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs.
7.3 Fine-tuning narrows the ICL gap between Mamba and Transformers
We next explore how MPFT and PEFT affect Mamba ICL performance. All Mamba pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples). We use the training recipe of  [50], which includes BF16 utilization. Performance is evaluated using the datasets from Table 1–HellaSwag [61], PIQA [6], Arc-E [9], Arc-C [9], and WinoGrande [46]–and report the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS). For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted)..
Figure 3 displays AIPSS for pretrained and instruction fine-tuned Mamba and Pythia models. As previously noted, pretrained Mamba models do not display similar ICL ability as comparable Pythia models on the evaluated standard NLP benchmarks. In particular, Mamba 2.8B, the largest pretrained Mamba model, displays inconsistent zero-shot improvements as the number of shots increase. However, after fine-tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase. Compared to Mamba pretrained models, which are only capable of 38% of the AIPSS compared to similar pretrained Pythia models, fine-tuned ALL LoRA Mamba models are capable of 81.5% of the AIPSS compared to similarly fine-tuned Pythia models.
(a) Mamba Pretrained
(b) Pythia Pretrained
(c) Mamba ALL LoRA
(d) Pythia ALL LoRA
Figure 3: Fine-tuning narrows the ICL gap between Mamba and Pythia. ALL LoRA models were instruction fine-tuned on the OpenHermes [49] dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard benchmarks.
7.3.1 Fine-tuning robustness
We show that Mamba is robust to the choice of PEFT hyperparemters. We conduct an extensive hyperparameter search across the learning rate, LoRA dimension, and number of warmup steps. From the Cartesian-product of these three parameters, 150 hyperparameter configurations were sampled and used to fine-tune Mamba 370M over the Openhermes dataset. For comparison, Pythia 410M is similarly fine-tuned using the same set of 150 hyperparameter configurations.
(a) Mamba 370M
(b) Pythia 410M
Figure 4: Fine-tuning hyperparameter search for OpenHermes. Each point is a different hyperparameter configuration. SLL LoRA was used for both models. The $x$-axis is the learning rate, the $y$-axis is resulting MMLU 5-shot performance, bubble size is the LoRA dimension, and the color is the number of warmup steps $\in\{0,1\mbox{k},2\mbox{k}\}$.
The MMLU 5-shot performance for each of the 150 Mamba and Pythia fine-tuned models is displayed in Figure 4. Pythia 410M is capable of higher performance than Mamba 370M, where the average accuracy for the former and the latter are 26.5% and 24.8%, respectively. However, Mamba 370M is much more robust to the choice of hyperparameters, with a difference of 1.5% between the minimum (23.3%) and maximum (24.8%). In contrast, Pythia 410M fine-tuned models display a large performance difference of 4.7% between the minimum (22.9%) and maximum (27.6%).
8 Discussion
We’ve extensively explored Mamba’s downstream learning capabilities. Using dynamical systems theory, we’ve shown that Mamba’s recurrent dynamics are robust to small input perturbations (contrary to the current understanding of Mamba’s recurrent sensitivities). We’ve extensively confirmed this result, showing that: a) Mamba inference is robust to changes due to mixed-precision, (b) Mamba inference differences due to mixed-precision align with Transformers, (c) Mamba fine-tuning is robust to changes due to mixed-precision and PEFT, and (d) differences in downstream performance for Mamba due to MPFT and PEFT can be more robust than Transformers. Using both MPFT and PEFT, we’ve shown that instruction fine-tuning Mamba SSMs greatly narrows the previously observed ICL gap, going from only 38% (post pretraining) up to 81.5% (post fine-tuning) of the ICL abilities of similar Transformers. Furthermore, we’ve shown that combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models.
There are significant avenues for future work. In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups. Furthermore, while the largest pure Mamba model contains 2.8B parameters, the training speedups and improved memory utilization described herein may be applied to more efficiently pretrain larger pure Mamba SSMs (e.g., 7B parameters and greater), where Mamba models may better manifest emergent abilities previously displayed by Transformers (or even manifest previously unobserved abilities).
8.1 Broader Impact
The Mamba models considered are all LLMs, and thus have the same potential positive and negative societal impacts as other LLMs (e.g., hallucinations). Furthermore, fine-tuning is known to possibly erode existing LLM guardrails, and thus our methods may be adapted for this fine-tuning use case (as is the case for all PEFT and MPFT methods). However, our work improves the quality of Mamba models for downstream applications, which may be adapted for all positive LLM applications in society (e.g., personal assistants, task automation, code completion, etc.). Finally, our work decreases the computational constraints required to train and inference Mamba SSMs, which has implications for green ML (e.g., decreased CO2 emissions, positive climate change impact, etc.). 410 GPU days were used to produce the results for this paper.
9 Acknowledgements
We thank Leidos for funding this research through the Office of Technology. Approved for public release 24-LEIDOS-0515-27826.
Appendix A Mamba stable dynamics proof
Recall the state-space parameters and equations for the MambaBlock; $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,n\}=[n]$. Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{n}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{n}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{n}\in\mathbb{R}{}^{d}$:
$\bm{x}_{t}$ $=\bar{\mathbf{A}}_{t}\bm{x}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{u}_{t}$ (5)
$\mathbf{y}_{t}$ $=\bar{\mathbf{C}}_{t}\bm{x}_{t},$
where $\bar{\bm{\Delta}}_{t}=\texttt{softplus}(\texttt{Linear}(\bm{\Delta}_{t}))\in\mathbb{R}_{\geq 0}{}^{d\times d}$, $\bar{\mathbf{A}}_{t}=\exp{(\bar{\bm{\Delta}}_{t}\mathbf{A})}$, $\bar{\mathbf{B}}_{t}=\mathbf{A}^{-1}(\bar{\mathbf{A}}-\mathbf{I})\mathbf{B}_{t}$, and $\mathbb{R}_{\geq 0}{}$ is the set of non-negative real numbers. In practice, $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are diagonal matrices.
Furthermore, recall the following definitions:
$\bm{x}_{t}$ $=F_{\theta}(\bm{x}_{t-1},\mathbf{u}_{t})$
where $\theta$ denotes the aforementioned time-varying parameters. For input sequence $\mathbf{u}_{t},\dots,\mathbf{u}_{T}$ and initial latent state value $\bm{x}_{0}$, we thus write
$\bm{x}_{T}$ $=F_{\theta}(F_{\theta}(\dots F_{\theta}(\bm{x}_{0},\mathbf{u}_{1})))\coloneq F_{\theta}^{T-1}(\bm{x}_{0},\mathbf{u}_{1}).$
We first prove that, given two scalar $\varepsilon$-close inputs to a MambaBlock, their deviations do not grow exponentially as the number of recurrences increases (Lemma 1). The main result in the paper is subsequently proved.
Lemma 1.
For input $(\bm{x}_{0},\mathbf{u}_{1})$ to a MambaBlock, small changes $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
Proof.
Firstly, we note that within the MambaBlock, $\mathbf{A}$ is stored in log-space followed by a negative exponentiation prior to use. Thus, $\mathbf{A}\in\mathbb{R}_{\leq 0}{}^{d\times d}$, where $\mathbb{R}_{\leq 0}{}$ is the set of non-positive real numbers.
Recall that for the maximum deviation, we have:
$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\lambda_{\mathtt{max}}{})}).$
where the maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ is defined as:
$\lambda_{\texttt{max}}$ $\coloneq\lim_{T\to\infty}\frac{1}{T}\log\left\lVert\prod_{t=0}^{T}\frac{\partial\bm{x}_{t}}{\partial\bm{x}_{t-1}}\right\rVert_{2},$
and $\left\lVert\right\rVert_{2}$ denotes the spectral norm for matrices.
Thus, to complete the proof, it suffices to show that $\lambda_{\mathtt{max}}{}\leq 0$. Recall that $\mathbf{A}$ and $\bar{\bm{\Delta}}_{t}$ are diagonal. From Equation 5, we thus have
$\lambda_{\texttt{max}}$ $=\lim_{T\to\infty}\frac{1}{T}\log\left\lVert\prod_{t=0}^{T}\frac{\partial\bm{x}_{t}}{\partial\bm{x}_{t-1}}\right\rVert_{2}$
$=\lim_{T\to\infty}\frac{1}{T}\log\left\lVert\prod_{t=0}^{T}\exp{(\bar{\bm{\Delta}}_{t}\mathbf{A})}\right\rVert_{2}$
$=\lim_{T\to\infty}\frac{1}{T}\log\left\lVert\exp{\sum_{t=0}^{T}(\bar{\bm{\Delta}}_{t}\mathbf{A})}\right\rVert_{2}$
Let $i$ be the dimension which corresponds to the output of the spectral norm, i.e., $i=\mbox{argmax}_{j=1,\dots,d}\{\exp{\sum_{t=0}^{T}(\bar{\bm{\Delta}}_{t}[j,j]\mathbf{A}[j,j])}\}$. We thus have
$\lambda_{\texttt{max}}$ $=\lim_{T\to\infty}\frac{1}{T}\log\left\lVert\exp{\sum_{t=0}^{T}(\bar{\bm{\Delta}}_{t}\mathbf{A})}\right\rVert_{2}$
$=\lim_{T\to\infty}\frac{1}{T}\log\exp{\sum_{t=0}^{T}(\bar{\bm{\Delta}}_{t}[i,i]\mathbf{A}[i,i])}$
$=\mathbf{A}[i,i]\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T}\bar{\bm{\Delta}}_{t}[i,i]$
$\mathbf{A}[i,i]$ is non-positive and $\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T}\bar{\bm{\Delta}}_{t}[i,i]\geq 0$, since $\bar{\bm{\Delta}}_{t}[i,i]\in\mathbb{R}_{\geq 0}{}\;\forall t$. Thus, $\lambda_{\mathtt{max}}\leq 0$. ∎
Theorem 1.
Let $(\bm{x}_{t-1},\mathbf{u}_{t})$ be the latent state and input at an arbitrary time $t\in[1,T]$ within a MambaBlock. Then small changes $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially decreasing over discrete-time, i.e., $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
Proof.
Let $\tau(t)$ be a function that maps time values such that $\tau(t)\in[1,T-t]$ and $\tau(t)=1,\tau(t+1)=2,\dots,\tau(t+T)=T-t$. Then $\mathbf{B}_{\tau(t)},\mathbf{C}_{\tau(t)},\bm{\Delta}_{\tau(t)}$ define a new MambaBlock with inputs $\mathbf{u}_{\tau(t)},\dots,\mathbf{u}_{\tau(t+T)}$ and subsequent recurrent states $\bm{x}_{\tau(t)},\dots,\bm{x}_{\tau(t+T)}$. Applying Lemma 1 to this MambaBlock with $(\bm{x}_{\tau(t)-1},\mathbf{u}_{\tau(t)})$ completes the proof. ∎
Appendix B Proof of weight-tying using LoRA in the MambaBlock
Theorem 2.
Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 3. Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
Proof.
Let $r$ be the specified LoRA dimension. Targeting this matrix for LoRA results in the adapter
$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
where $\mathbf{U}\in\mathbb{R}{}^{n\times r}$, $\mathbf{V}\in\mathbb{R}{}^{r\times 3d}$, and $\mathbf{W}$ is frozen during fine-tuning. Thus, for index $[i,j]$,
$\mathbf{W}^{\prime}[i,j]=$ $\sum_{k=0}^{r-1}\mathbf{U}[i,k]\mathbf{V}[k,j].$
Recall the form of $\mathbf{W}$:
$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$
where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. For index $[t-1,j]$, we thus have
$\tilde{\mathbf{W}}[t-1,j]=$ $\mathbf{W}[t-1,j]+\mathbf{W}^{\prime}[t-1,j]$
$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
Thus, the weights $\mathbf{U}[t-1,:]$ are tied for any parameter $\tilde{\mathbf{W}}[t-1,j],j\in\{1,\dots,3d\}$, which are used to adapt parameters $\bm{\Delta}_{1},\mathbf{B}_{t},$ and $\mathbf{C}_{t}$.
Appendix C Experimental Details
All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16], version 0.4.2. Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models. All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0.
For MPFT, Flash Attention 2.0 [10] via flash_attn 2.5.7 was used for Pythia models. For FP16 and BF16 inference results, Flash Attention 2.0 was used for both Pythia and OLMo models. For OLMo results, the 336B-token checkpoint was used by specifying revision=step80000-tokens336B.
Outside of the OpenHermes hyperparameter search, all Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps. Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively. For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 3).
For SLL LoRA, targeted Mamba layers were {x_proj, embeddings, in_proj, out_proj}; x_proj is the large MambaBlock memory buffer which, when targeted by LoRA, regularizes the majority of SSM parameters during fine-tuning through weight tying (Theorem 2). Pythia targeted SLL LoRA layers were {dense, embed_in, query_key_value, dense_h_to_4h,dense_4h_to_h}, chosen to balance performance across model sizes.
All experiments in Tables 1 and  2, Figures 1 and  2(b) were run using a single-GPU Nvidia A10G (24 GB total memory). For Pythia and Mamba ALL LoRA experiments in Figure 3, all experiments were run on an A10G, except for Mamba 2.8B, which exceeded A10G memory capacity and was run on an Nvidia H100 (80 GB total memory).
Table 3: Learning rate and LoRA dimension $r$ values
Mamba size Pythia size learning rate LoRA $r$
130M 160M 1.0e-5 8
370M 410M 5.0e-5 16
790M 1B 1.0e-6 32
1.4B 1.4B 5.0e-6 64
2.8B 2.8B 5.0e-7 128
For the hyperparameter search results in Figure 4, all experiments were run using 8 H100 GPUs. SLL LoRA was used for Mamba and Pythia models. The range of hyperparameter values was as follows:
  • learning rate $\in\{1e-7,2e-7,5e-7,1e-6,2e-6,5e-6,1e-5,2e-5,5e-5,1e-4,2e-4,5e-4,1e-3,2e-3,5e-3\}$
  • LoRA dimension $r\in\{16,32,64,128,256\}$
  • warmup steps $\in\{0,1000,2000\}$
All other hyperparameters followed previous experiments.
The Alpaca dataset is freely available for download at https://huggingface.co/datasets/tatsu-lab/alpaca under open-source license CC-by-NC 4.0. The OpenHermes dataset is freely available for download at https://huggingface.co/datasets/teknium/OpenHermes-2.5 under open-source license MIT, Apache 2.0, CC.

### v2

@@
-Mamba State-Space Models Can Be Strong Downstream Learners
+Mamba State-Space Models Are Lyapunov-Stable Learners
 Abstract
-Mamba [22] state-space models (SSMs) have recently outperformed state-of-the-art (SOTA) Transformer large language models (LLMs) in various tasks and been widely adapted.
-However, Mamba’s downstream learning capabilities remain either unexplored–e.g., mixed-precision (MPFT) and parameter-efficient fine-tuning (PEFT)–or under-evaluated–e.g., in-context learning (ICL).
-For the latter, recent works [42, 19] reported Mamba’s ICL rivals SOTA Transformer LLMs using non-standard benchmarks.
-In contrast, we show that on standard benchmarks, pretrained Mamba models achieve only 38% of the ICL performance improvements (over zero-shot) of comparable Transformers.
-Enabling MPFT and PEFT in Mamba architectures is challenging due to recurrent dynamics and highly customized CUDA kernels, respectively.
-However, we prove that Mamba’s recurrent dynamics are robust to small input changes using dynamical systems theory.
-Empirically, we show that performance changes in Mamba’s inference and fine-tuning due to mixed-precision align with Transformer LLMs.
-Furthermore, we show that targeting key memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus achieving parameter efficiency while retaining speedups.
-We show that combining MPFT and PEFT enables up to 2.15 times more tokens-per-second and 65.5% reduced per-token-memory compared to full Mamba fine-tuning, while achieving up to 81.5% of the ICL performance improvements (over zero-shot) of comparably fine-tuned Transformers.
-**footnotetext: Equal Contribution. Correspondence to john.t.halloran@leidos.com.
+Mamba [21, 11] state-space models (SSMs) were recently shown to outperform state-of-the-art (SOTA) Transformer large language models (LLMs) across various tasks.
+Despite subsequent widespread adaptation, little work has focused on Mamba LLMs’ amenability for fine-tuning frameworks ubiquitously used for Transformer-based LLMs, e.g., mixed-precision fine-tuning (MPFT) and parameter-efficient fine-tuning (PEFT).
+For the former, it currently remains an open question whether Mamba’s recurrent dynamics are robust to small input changes, such as those encountered during MPFT.
+Using dynamical systems theory (in particular, Lyapunov exponents), we answer this question in the affirmative.
+We empirically validate this result through several experiments, showing that Mamba SSMs are significantly more stable to changes introduced by mixed-precision than comparable Transformers, even when both MPFT and PEFT are combined.
+For PEFT, we show how targeting specific memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus providing both parameter efficient learning and computational savings.
+Finally, with both MPFT and PEFT enabled, we explore the impact of instruction tuning Mamba SSMs for in-context learning (ICL) on natural language tasks.
+While pretrained Mamba and Mamba-2 models only achieve 38% and 82% (respectively) of the ICL improvements of comparable Transformer-based LLMs, we show that instruction tuning allows Mamba models to narrow this gap to 81% and Mamba-2 models to skyrocket over this gap to 132%.
+**footnotetext: Correspondence to john.t.halloran@leidos.com.
 1 Introduction
-Innovating on previous state-space models (SSMs) [23, 11], Mamba [22] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs).
-Mamba was initially shown to greatly outperform comparable Transformer LLMs [5] across a large number of standard natural language benchmarks.
-Subsequently, pretrained Mamba models have been widely adapted across different data modalities [39, 62, 34, 43, 35], tasks [57, 59, 45, 60, 54, 35, 2], and architectures [1, 42, 37].
-However, despite such rapid and widespread adaptation, evaluation of Mamba’s ability to perform standard downstream learning abilities exhibited by Transformer-based LLMs have either not been extensively conducted on standard natural benchmarks or are completely lacking.
-For instance, while recent works [42, 19, 29] have evaluated Mamba’s ability to perform in-context learning (ICL), such studies focused extensively on either non-natural tasks [29, 17] or non-standard benchmarks [25].
-Furthermore, evaluation of Mamba’s mixed-precision fine-tuning (MPFT) and performance efficient fine-tuning (PEFT) capabilities are currently lacking.
-For the former, MPFT (and, by extension, mixed-precision inference) are made difficult due to potential sensitivities of Mamba’s recurrent dynamics, where  [21, 28] suggest full precision (FP32) is required to perform stable training.
-For the latter, PEFT via standard low-rank adaptation (LoRA) [27] is made difficult within Mamba’s SSM layer (referred to herein as the MambaBlock) due to highly customized SSM CUDA kernels which provide competitive performance to attention-based speedups [10] at the cost of standard adapter support.
-However, PEFT and MPFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [52], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12].
-Herein, we extensively explore Mamba’s downstream learning capabilities across standard natural benchmarks.
-For ICL, we show that, in contrast to recent non-standard studies showing Mamba models rival state-of-the-art (SOTA) LLMs of similar parameter counts, the pretrained benefits of Mamba few-shot learning are significantly less than comparable Transformer LLMs across standard natural benchmarks; averaged across the benchmarks and parameter counts in Table 1, Mamba models only achieve 38% of the performance improvements (relative to zero-shot) of comparable Transformer models from the Pythia suite [5].
-However, we show in the sequel that Mamba models can more than halve this gap through efficient fine-tuning, achieving as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers.
-For MPFT, we leverage theory from dynamical systems to show that small input changes in a MambaBlock do not lead to exponentially deviating outputs.
-Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference and fine-tuning are on par with those demonstrated by Transformer LLMs (Section 7).
-For PEFT, we show that by targeting the largest memory buffer exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying.
+Innovating on previous state-space models (SSMs) [22, 10], Mamba [21, 11] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs).
+Upon their introduction in [21], Mamba SSMs were shown to greatly outperform comparable attention-based LLMs [4] across a large number of standard natural language benchmarks.
+Subsequently, pretrained Mamba models have been widely adapted across a large number of data modalities [36, 31, 43, 32], tasks [57, 53], and architectures  [1, 40, 34].
+Despite such widespread adaptation and subsequent research threads [11, 40, 54], little work has been done to understand the amenability of Mamba SSMs for widely used fine-tuning frameworks, such as mixed-precision fine-tuning (MPFT) [38] and parameter-efficient fine-tuning (PEFT) [23, 25].
+MPFT and PEFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [51], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12].
+However, direct application of MPFT for Mamba SSMs is made difficult due to potential sensitivities of Mamba’s state-space dynamics, a common concern for recurrent-based deep models [41].
+To combat this, both [26] and [20] suggest full precision (FP32) may be required to perform stable training for Mamba models.
+Thus, it is currently an open question whether Mamba’s recurrent dynamics are stable in the presence of small input deviations, such as those introduced in MPFT.
+To answer this question, we leverage theory from dynamical systems.
+Deriving and bounding the Lyapunov exponents for both Mamba and Mamba-2 models, we show that small input changes within the SSM layer of either model do not lead to exponentially deviating outputs.
+Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference are on par with those demonstrated by Transformer LLMs, while deviations due to MPFT are significantly more stable than those of comparable Transformers (Section 5).
+Furthermore, this trend continues when MPFT and PEFT are combined, where Mamba SSMs again produce significantly smaller deviations compared to comparable Transformer LLMs.
+For PEFT, we show that by targeting the large memory buffers exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying.
 We show that this leads to extremely efficient PEFT, resulting in up to 2.15 times faster training and 65.5% reduced memory compared to the largest evaluated Mamba model without MPFT or PEFT.
-Table 1: In-context learning performance for pretrained Mamba and Pythia models.
-Models are collected into parameter classes for head-to-head comparison using the groupings in [22].
-Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16].
-LAMBADA zero-shot is more effective for the model sizes considered (further discussed in [58, 8]) and thus excluded from few-shot performance averages.
-Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
-Model $N$-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
-ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
-Mamba 130M 0 16.07 44.3 35.3 64.5 48.0 24.2 44.8
-1 19.34 38.3 35.2 64.3 47.1 23.5 51.3 -1.4
-3 23.13 35.4 35.1 65.1 49.0 24.0 50.7 -0.2
-5 24.38 36.2 34.8 64.9 49.2 23.9 50.5 -0.5
-[…]
-1 47.21 28.2 30.6 62.2 43.4 23.7 49.3 -0.4
-3 63.70 24.7 30.5 61.9 44.8 22.9 51.3 0.1
-5 66.30 25.3 30.4 62.6 43.4 23.1 50.8 -0.2
-Mamba 370M 0 8.14 55.6 46.5 69.5 55.0 27.9 55.5
-1 9.74 49.8 45.9 69.3 57.4 26.5 54.6 -0.8
-3 10.89 48.5 46.2 69.6 58.7 28.5 53.6 1.0
-5 11.36 48.5 46.2 69.4 58.3 28.0 56.0 1.3
-[…]
-1 12.26 47.1 40.5 68.0 53.8 25.6 52.4 1.8
-3 14.39 43.2 40.9 67.9 55.1 26.9 54.0 4.2
-5 14.62 44.1 40.8 68.1 54.6 26.6 53.4 3.5
-Mamba 790M 0 6.01 61.7 55.1 72.1 61.2 29.6 56.0
-1 7.06 56.2 54.5 72.5 63.3 30.1 56.9 1.4
-3 8.05 54.8 54.2 72.2 63.4 31.6 57.2 2.4
-5 8.83 53.4 54.6 72.5 64.6 32.1 57.5 3.4
-[…]
-1 8.99 51.8 47.3 70.7 57.1 28.2 53.4 1.0
-3 10.48 48.2 47.5 71.2 59.2 28.0 54.3 2.2
-5 10.86 48.4 47.3 71.4 58.7 28.4 53.1 1.9
-Mamba 1.4B 0 5.04 65.0 59.1 74.2 65.5 32.9 58.6
-1 5.83 60.6 58.20 74.7 64.5 33.0 61.2 -0.5
-3 6.62 58.9 58.8 73.7 66.1 34.4 60.9 0.6
-5 6.98 58.4 59.0 74.0 66.4 35.5 60.5 1.4
-[…]
-1 6.96 56.3 52.1 71.4 62.0 29.5 57.5 1.4
-3 7.89 54.4 52.6 70.9 63.9 31.1 56.8 2.9
-5 8.02 54.4 52.8 71.0 63.2 31.3 57.8 3.3
-Mamba 2.8B 0 4.23 69.2 66.2 75.2 69.7 36.3 63.4
-1 5.01 63.9 65.7 75.5 69.8 37.2 63.7 0.6
-3 5.53 63.0 65.5 75.2 70.8 38.1 64.8 1.6
-5 5.70 62.7 66.2 76.2 70.9 38.3 64.6 2.1
-[…]
-1 5.66 60.9 59.4 73.8 66.8 34.8 59.0 1.7
-3 6.20 59.1 59.9 74.7 67.4 34.9 60.8 2.9
-5 6.52 59.1 60.2 74.5 67.1 35.0 61.3 3.1
+Furthermore, this allows even the largest (2.8 billion parameter) Mamba LLMs to be fine-tuned on a single GPU with as little as 24GB of onboard memory.
+Finally, using both MPFT and PEFT, we complement existing studies [40, 30] by exploring the ICL capabilities of instruction-tuned Mamba and Mamba-2 models on natural language tasks.
+In particular, the ICL capabilities of both pretrained Mamba and Mamba-2 models lag behind those of comparable Transformer models from the Pythia suite [4]; averaged across five standard natural language benchmarks and foundation model sizes, Mamba and Mamba-2 models only achieve 38% and 82%, respectively, of the performance improvements (relative to zero-shot) of Pythia models.
+However, after instruction-tuning, Mamba models are able to achieve as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers, while Mamba-2 models push this to 132% of the ICL improvements achieved by Pythia models.
+We note that, similar to Transformer foundation models [55], (post) instruction tuning ICL appears as an emergent abilities for Mamba and Mamba-2 SSMs, manifesting for models of size 370 million parameters and greater, while failing to manifest for Mamba and Mamba-2 models of fewer parameter counts.
+We note that herein, when describing a particular foundation model or result, we use the term “Mamba model” to refer to one of the original models released in [21] and “Mamba-2 model” to refer to models released in [11].
+While there are subtle architectural differences between these two SSMs, they share important similarities which allow our theoretical results to extend to both sets of models.
+In particular, Mamba and Mamba-2 models share the same state-space equations, support for SSM matrices, and design scheme of storing the majority of SSM parameters in a large memory buffer.
+Thus, we synonymously use the term MambaBlock to refer to the SSM layer of both Mamba and Mamba-2 models.
 2 Background
-2.1 Downstream learning for LLMs
-Since the release of the Transformer architecture [51], attention-based LLMs have exhibited several downstream learning abilities–in particular, PEFT, MPFT, and ICL–which allow the rapid adaptation of foundation models towards specific applications.
+Mixed-precision and parameter-efficient fine-tuning.
+In the current era of extremely large foundation models, both MPFT and PEFT have become ubiquitous tools for the rapid adaptation of Transformer-based LLMs towards specific applications.
 PEFT using adapters [23] allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters.
@@
 We note that post-fine-tuning quantization approaches [13] may be further used to decrease Transformer LLM computational demands, but such approaches are not considered in this work.
-ICL provides an adaptable alternative to fine-tuning.
-Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest.
-Given sufficiently large models and pretraining data [8, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting.
-ICL thus avoids computational and storage costs associated with fine-tuning and PEFT adapter weights, respectively, and rather changes the prompt to immediately change the model for downstream learning.
-However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [8, 38].
-2.2 State-space Models
+State-space Models.
 Structured state-space sequence (S4) models [22, 14] are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency.
@@
 To enable hardware efficiency with time-varying SSM parameters, [21] thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states.
+Subsequently, [11] considered the unrolled state-space equations and leveraged tensor contractions (i.e., einsum notation [45]) to efficiently calculate Mamba variables.
+The resulting Mamba-2 foundation models contained significantly larger latent-variable dimensions than the Mamba models of [21], while maintaining efficiency on modern GPU accelerators.
+In-context learning.
+ICL provides an adaptable alternative to fine-tuning.
+Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest.
+Given sufficiently large models and pretraining data [7, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting.
+However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [7, 35].
 3 Mamba state-space models
-For model dimension $d$ and maximum input sequence length $T$, the MambaBlock defines state-space parameters $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}$.
+For latent-variable dimension $d$ and maximum input sequence length $T$, the MambaBlock defines state-space parameters $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}$.
 The matrix $\bm{\Delta}_{t}$ controls the discrete step-size.
@@
 In practice, $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are diagonal matrices.
-3.1 Hardware-aware optimizations
-As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [11] are no longer applicable.
-However, by diagonality, each dimension may be computed in parallel.
-Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [7].
-[15] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated.
-Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{n\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
-$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$ (3)
-where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on.
-4 Stable dynamics in the MambaBlock
+3.1 Stable dynamics in the MambaBlock
 The Mamba foundation models were pretrained in full FP32 precision.
@@
 That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
-The proof of Theorem 1 is available in Appendix A, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
-4.1 Consequences for automatic mixed-precision
+The proof of Theorem 1 is available in Appendix B, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
+Thus, the latent states of Mamba and Mamba-2 models are stable under small input changes.
+However, variables $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ are the primary outputs for such models, particularly for LLM applications.
+We next show that, given Theorem 1, Mamba and Mamba-2 output variables are also stable.
+Theorem 2.
+Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time.
+Then small changes to the output $\mathbf{y}_{t}$ are also exponentially non-increasing over discrete time.
+The proof of Theorem 2 is available in Appendix C. Thus, by Theorems 1 and  2, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during recurrency.
+3.1.1 Consequences for automatic mixed-precision
 During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16).
 During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating.
-In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ fed into the SSM scan (which are passed through a SwiGLU), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
+In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
 For a discrete dynamical system with $\lambda_{\texttt{max}}>0$, changes due to AMP compound after repeated expansion of the recurrent state, thus leading to exponential deviations between quantities calculated using mixed- versus full-precision.
@@
 However, by Theorem 1, such differences do not exponentially compound over discrete-time within the MambaBlock.
-5 Regularized fine-tuning using LoRA in the MambaBlock
-Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer.
-E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these invidiual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer).
-Rather than adjusting Mamba’s low-level CUDA kernels themselves to integrate LoRA within the highly optimized prefix scan, we note that $\mathbf{W}$ can instead be directly targeted.
-Doing so, we have the following, where the proof is available in Appendix B.
+3.2 Hardware-aware optimizations and PEFT
+As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [10] are no longer applicable.
+However, by diagonality, each dimension may be computed in parallel.
+Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [6].
+[21] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated.
+Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{T\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
+$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$ (4)
+where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on.
+The customized Mamba prefix scan kernel heavily relies on this memory layout to optimize the access pattern of $\mathbf{W}$ in Equations 1 and 2.
+Similarly, Mamba-2 stores diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ in a large memory buffer $\mathbf{W}$.
+However, rather than utilizing the underlying recurrence to directly compute the hidden state and output at each time-step, [11] consider the matrix resulting from unrolling Equation 2 across all $t$.
+Mamba-2 models thus leverage the structure of the ensuing semiseparable matrix to calculate $\bm{x}_{1},\dots,\bm{x}_{T}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ using tensor contractions, which are highly optimized on modern hardware accelerators.
+When fine-tuning using LoRA, low-rank matrices are used to adapt the frozen weight matrices of targeted linear layers.
+The importance of $\mathbf{W}$ for both Mamba and Mamba-2 models makes it a primary candidate for LoRA adaptation.
+In such cases, selecting $\mathbf{W}$ for LoRA adaptation results in the following:
 Theorem 3.
@@
 Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
-Thus, the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
+The proof of Theorem 3 is available in Appendix A. The specific affects of both targeting $\mathbf{W}$ for fine-tuning and Theorem 3’s impact on generalization are ablated in Appendix F.
 4 Related Work
-Several recent works [42, 19, 29, 37] have studied Mamba’s ability to perform ICL.
-However, none of these have extensively studied Mamba’s ICL capabilities either on standard NLP benchmarks or on pure MambaBlock foundation models.
-In particular, foundational Mamba models’ ICL abilities were tested in [42] to learn simple function classes (e.g., logistic regression and decision trees [17]) and in  [19] to learn non-standard NLP benchmarks (i.e., task vectors [25]).
-While [42, 19] report Mamba’s ICL abilities rival SOTA Transformers, their utilized benchmarks were proposed as supplemental ICL studies after Transformer LLMs’ success on standard NLP benchmarks [8].
-Indeed, direct evaluation of Mamba foundation models on standard NLP benchmarks does not lead to higher gains over zero-shot performance relative to comparable Transformer LLMs (demonstrated in Table 1).
+Recent work has sought to understand how to efficiently increase Mamba’s hidden-state dimension by restructuring SSM operations using tensor contractions [11], leading to Mamba-2 models.
+A separate line of work has sought to understand the in-context learning (ICL) capabilities of Mamba LLMs when trained from scratch for specific tasks [40, 30].
+Another line of recent work has sought to understand how hybrid Mamba-Transformer models may be directly distilled from Transformer models [54].
+However, to the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or sought to understand the effects of LoRA adaptation on modules within the MambaBlock.
 Lyapunov exponents have previously been considered for classic RNN structures (e.g., vanilla RNNs, LSTMs, GRUs, PLRNNs, etc.) [39, 52], to determine when such models exhibit chaotic dynamics and the impact on the exploding/vanishing gradient phenomena***We note that this continues a long line of research exploring RNNs sensitivity to initial conditions and their subsequent ability to produce chaotic output [44, 29, 2, 3], although previous work did not leverage Lyapunov exponents..
@@
 To the best of our knowledge, no previous works have used Lyapunov exponents to explore the effects of mixed-precision on recurrent neural models or Mamba architectures.
-As in  [22], the majority of subsequent Mamba works have focused on pretraining MambaBlocks using full precision [62, 59, 1, 37].
-Notably, the official implementation of Jamba [37], the Transformer-Mamba hybrid, supports mixed- and 8-bit precision, but avoids MambaBlocks when applying such quantization [31].
-Similarly, the official Mamba sources advise using full precision within the MambaBlock [28, 21], cautioning against using mixed-precision due to potential recurrent sensitivities.
-To the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or explored pure Mamba networks fine-tuning abilities relative to Transformer LLMs.
+As previously noted, recent works [40, 30] have studied Mamba’s ability to perform ICL by training Mamba models for specific tasks.
+Such tasks include logistic regression, decision trees, and learning other simple function classes, following the work of [17].
+We emphasize that, in this set up, relatively small Mamba models–33 million and 90 million parameters for [30] and [40], respectively–are trained from scratch for every evaluated task.
+Indeed, [40] notes that subsequent work is necessary to understand Mamba’s ICL capabilities for language modeling using standard natural language benchmarks, as well as for larger model sizes.
+Thus, our study of both the pretrained and instruction tuned ICL capabilities of Mamba/Mamba-2 LLMs for natural language tasks are complimentary to previous works.
 5 Experiments
 To demonstrate the implications of Theorem 2, we explore the performance difference between running inference with full-precision pretrained weights and using mixed-precision (FP16 and BF16) weights.
-Model performance is measured as percent accuracy using the MMLU [26] dataset.
+Model performance is measured as percent accuracy using the MMLU dataset [24].
 The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy.
 Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy.
+Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [4], OLMo [19] 336B-token checkpoint, and Phi 1.5 [33]).
+We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers.
+To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory).
+All models were evaluated using the LM evaluation harness from Eleuther AI [16].
+Further experimental details are available in Appendix D.
+The results are available in Table 1.
 Table 1: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference.
@@
 BF16 $\mu$ 0.05 1.45 0.20 0.20 0.66 0.16 0.13 0.31 0.13 1.05 1.17 0.11
-Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [5], OLMo [20] 336B-token checkpoint, and Phi 1.5 [36]).
-We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers.
-To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory).
-All models were evaluated using the LM evaluation harness from Eleuther AI [16].
-Further experimental details are available in Appendix C.
-The results are available in Table 2.
 From Table 1, inferencing in Pythia using FP16 and BF16 result in an average 0.13 and 0.41 full-precision divergence, respectively.
@@
 Interestingly, both SSM and Transformer architectures exhibit large divergence spikes–i.e., mean divergence greater than a percentage point–when using BF16, which occurs once for Mamba and Phi 1.5 models and twice for Pythia models.
-In the following, we show that such spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
-7.1 Non-divergent Mamba fine-tuning
+Due to space constraints, Mamba-2 results for the same experiment are included in Appendix Table 3.
+We note that, for comparable model sizes and mixed-precision, Mamba-2 models follow identical divergence trends as respective Mamba models.
+In the following, we show that the observed large divergence spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
+5.1 Mamba SSMs are stable learners
 We next explore the implications of Theorem 1 on fine-tuning, wherein mixed-precision is especially critical; MPFT combined with PEFT adapters have been shown to drastically reduce Transformer fine-tuning times [12].
@@
 We focus on utilizing LoRA [25], which is arguably the most widely used PEFT framework for LLMs.
+Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba, Pythia, and OpenELM models.
+Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT.
+Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
 Using the Alpaca dataset [48], Mamba 160M, 410M, and 790M models are fine-tuned for three epochs with a maximum sequence length of 512.
@@
 While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules.
-For comparison, Pythia 160M, 410M, and 1B models are fine-tuned using the same experimental setup.
+For comparison, two Transformer-based LLM families of similar parameter counts are fine-tuned using the same experimental setup: Pythia (sizes 160M, 410M, and 1B) and OpenELM [37] (sizes 270M and 450M).
 The training recipe for all models was adapted from [50], with the AdamW_torch optimizer and a cosine annealing schedule.
-Further experimental details are available in Appendix C.
-Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba and Pythia models.
-Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT.
-Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
-For each Mamba and Pythia model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy.
-Across mixed-precisions and adapter settings, Mamba displays comparable divergences to Pythia models.
-E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia.
-Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia.
-Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning (in contrast to Pythia models).
-7.2 Hardware throughput and memory-utilization improvements given PEFT and MPFT
-With comparable divergences to Transformers and stable dynamics, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs.
+Further experimental details are available in Appendix D.
+For each Mamba, Pythia, and OpenELM model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy.
+Across mixed-precisions and adapter settings, Mamba displays smaller divergences than both Pythia and OpenELM models.
+E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia and 0.54 for OpenELM.
+Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia and 0.33 for OpenELM.
+Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning, in contrast to both Pythia and OpenELM models.
+(a) Average tokens-per-second $\uparrow$.
+(b) Maximum memory-per-token $\downarrow$.
+Figure 2: Timing and memory usage calculated Mamba model-sizes and PEFT combinations.
+Each model was trained using the Alpaca dataset dataset for three epochs and maximum sequence length 512.
+For each PEFT combination, the batch size was tuned to maximize GPU occupancy.
+Full fine-tuning exceeds available GPU memory (24 GB) for models greater than 790 million parameters.
+Hardware throughput and memory-utilization improvements.
+With stable dynamics and observed divergences smaller than comparable Transformers, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs.
 To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset.
@@
 Results are plotted in Figure 2(b).
-(a) Average tokens-per-second $\uparrow$.
-(b) Maximum memory-per-token $\downarrow$.
-Figure 2: Timing and memory usage calculated Mamba model-sizes and PEFT combinations.
-Each model was trained using the Alpaca dataset [48] dataset for three epochs and maximum sequence length 512.
-For each PEFT combination, the batch size was tuned to maximize GPU occupancy.
 Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 2(b).
@@
 Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs.
-7.3 Fine-tuning narrows the ICL gap between Mamba and Transformers
-We next explore how MPFT and PEFT affect Mamba ICL performance.
-All Mamba pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples).
+Furthermore, while full fine-tuning is no longer possible on a single A10G for Mamba models greater than 790 million parameters, MPFT and PEFT allow training Mamba models up to 2.8 billion parameters on GPUs with as little as 24 GB onboard memory.
+5.2 Instruction tuning impact on Mamba ICL for natural language tasks
+Using both MPFT and PEFT, we next explore how instruction tuning affects Mamba and Mamba-2 ICL performance on natural language tasks.
+All Mamba and Mamba-2 pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples).
 We use the training recipe of  [50], which includes BF16 utilization.
-Performance is evaluated using the datasets from Table 1–HellaSwag [61], PIQA [6], Arc-E [9], Arc-C [9], and WinoGrande [46]–and report the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS).
-For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted)..
+Zero and few-shot performance is evaluated using five standard natural language benchmarks: HellaSwag [58], PIQA [5], Arc-E [8], Arc-C [8], and WinoGrande [46].
+ICL performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS).
+For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted).
+(a) Mamba Pretrained
+(b) Mamba-2 Pretrained
+(c) Pythia Pretrained
+(d) Mamba ALL LoRA
+(e) Mamba2 ALL LoRA
+(f) Pythia ALL LoRA
+Figure 3: Instruction tuning narrows the ICL gap between Mamba and Pythia, and creates a gap from Pythia to Mamba-2 models.
+ALL LoRA models were instruction tuned on the OpenHermes [49] dataset for one epoch.
+Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks.
 Figure 3 displays AIPSS for pretrained and instruction fine-tuned Mamba and Pythia models.
@@
 In particular, Mamba 2.8B, the largest pretrained Mamba model, displays inconsistent zero-shot improvements as the number of shots increase.
-However, after fine-tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase.
-Compared to Mamba pretrained models, which are only capable of 38% of the AIPSS compared to similar pretrained Pythia models, fine-tuned ALL LoRA Mamba models are capable of 81.5% of the AIPSS compared to similarly fine-tuned Pythia models.
-(a) Mamba Pretrained
-(b) Pythia Pretrained
-(c) Mamba ALL LoRA
-(d) Pythia ALL LoRA
-Figure 3: Fine-tuning narrows the ICL gap between Mamba and Pythia.
-ALL LoRA models were instruction fine-tuned on the OpenHermes [49] dataset for one epoch.
-Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard benchmarks.
-7.3.1 Fine-tuning robustness
-We show that Mamba is robust to the choice of PEFT hyperparemters.
-We conduct an extensive hyperparameter search across the learning rate, LoRA dimension, and number of warmup steps.
-From the Cartesian-product of these three parameters, 150 hyperparameter configurations were sampled and used to fine-tune Mamba 370M over the Openhermes dataset.
-For comparison, Pythia 410M is similarly fine-tuned using the same set of 150 hyperparameter configurations.
-(a) Mamba 370M
-(b) Pythia 410M
-Figure 4: Fine-tuning hyperparameter search for OpenHermes.
-Each point is a different hyperparameter configuration.
-SLL LoRA was used for both models.
-The $x$-axis is the learning rate, the $y$-axis is resulting MMLU 5-shot performance, bubble size is the LoRA dimension, and the color is the number of warmup steps $\in\{0,1\mbox{k},2\mbox{k}\}$.
-The MMLU 5-shot performance for each of the 150 Mamba and Pythia fine-tuned models is displayed in Figure 4.
-Pythia 410M is capable of higher performance than Mamba 370M, where the average accuracy for the former and the latter are 26.5% and 24.8%, respectively.
-However, Mamba 370M is much more robust to the choice of hyperparameters, with a difference of 1.5% between the minimum (23.3%) and maximum (24.8%).
-In contrast, Pythia 410M fine-tuned models display a large performance difference of 4.7% between the minimum (22.9%) and maximum (27.6%).
-8 Discussion
-We’ve extensively explored Mamba’s downstream learning capabilities.
-Using dynamical systems theory, we’ve shown that Mamba’s recurrent dynamics are robust to small input perturbations (contrary to the current understanding of Mamba’s recurrent sensitivities).
-We’ve extensively confirmed this result, showing that: a) Mamba inference is robust to changes due to mixed-precision, (b) Mamba inference differences due to mixed-precision align with Transformers, (c) Mamba fine-tuning is robust to changes due to mixed-precision and PEFT, and (d) differences in downstream performance for Mamba due to MPFT and PEFT can be more robust than Transformers.
-Using both MPFT and PEFT, we’ve shown that instruction fine-tuning Mamba SSMs greatly narrows the previously observed ICL gap, going from only 38% (post pretraining) up to 81.5% (post fine-tuning) of the ICL abilities of similar Transformers.
-Furthermore, we’ve shown that combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models.
-There are significant avenues for future work.
-In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups.
-Furthermore, while the largest pure Mamba model contains 2.8B parameters, the training speedups and improved memory utilization described herein may be applied to more efficiently pretrain larger pure Mamba SSMs (e.g., 7B parameters and greater), where Mamba models may better manifest emergent abilities previously displayed by Transformers (or even manifest previously unobserved abilities).
-8.1 Broader Impact
-The Mamba models considered are all LLMs, and thus have the same potential positive and negative societal impacts as other LLMs (e.g., hallucinations).
-Furthermore, fine-tuning is known to possibly erode existing LLM guardrails, and thus our methods may be adapted for this fine-tuning use case (as is the case for all PEFT and MPFT methods).
-However, our work improves the quality of Mamba models for downstream applications, which may be adapted for all positive LLM applications in society (e.g., personal assistants, task automation, code completion, etc.).
-Finally, our work decreases the computational constraints required to train and inference Mamba SSMs, which has implications for green ML (e.g., decreased CO2 emissions, positive climate change impact, etc.).
-410 GPU days were used to produce the results for this paper.
-9 Acknowledgements
+While pretrained Mamba-2 models display significantly better ICL ability than Mamba models, Mamba-2 models smaller than 780 million parameters struggle.
+However, after instruction tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase.
+Similarly, the majority of Mamba-2 models larger than Mamba 130M greatly improve in ICL performance.
+Thus, while pretrained Mamba and Mamba-2 models are only capable of 38% and 82% (respectively) of the AIPSS compared to similar pretrained Pythia models, instruction tuned Mamba and Mamba-2 models are capable of 81.5% and 133% of the AIPSS relative to similarly fine-tuned Pythia models.
+We note that a significant difference between Mamba and Mamba-2 models is the larger (by a factor of four) latent dimension.
+5.3 ICL is an emergent ability of instruction tuned Mamba SSMs
+We next study the emergent behavior (as a function of model size) of Mamba/Mamba-2 SSMs’ ICL abilities on natural language tasks by comparing to a larger number of Transformer-based LLMs of varying sizes.
+We compare to OpenELM [37] (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B [59], and OLMo 1.2B [19].
+To limit the emergent effects on both parameter size and pretraining token counts, we did not evaluate models greater than 2.8 billion parameters and chose open-source checkpoints as close as possible to the 300 billion total pretraining tokens used for Mamba, Mamba-2, and Pythia models.
+Thus, pretraining token counts for OpenELM, TinyLlama, and OLMo models were 429 billion, 503 billion, and 336 billion, respectively.
+We note this potentially biases ICL performance in favor of the newly evaluated Transformer-based LLMs, and that direct comparisons between Mamba, Mamba-2, and Pythia are the most fair (as these three classes of models were all pretrained on the same dataset for the same number of total pretraining tokens).
+We repeat the experiments from Figure 3, where we evaluate the pretrained and instruction tuned ICL capabilities of all models.
+To understand the critical role of parameter counts, we group all models into two classes: LLMs containing 450 million parameters or less, and LLMs containing greater than 450 million parameters.
+ICL performance measured by AIPSS is displayed in Figure 4.
+[…]
+[…]
+[…]
+[…]
+Figure 4: Instruction tuning improves Mamba-2 ICL performance past Transformer LLMs.
+ALL LoRA models were instruction fine-tuned on the OpenHermes dataset for one epoch.
+Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks: HellaSwag, PIQA, Arc-E, Arc-C, and WinoGrande.
+From Figure 4, it is clear that pretrained SSMs and Transformers of parameter counts 270 million and less display slight or detrimental ICL abilities (i.e., few-shot performance is worse than zero-shot).
+For models of greater than 450 million parameters, the majority of SSMs and Transformers display positive ICL abilities, with Mamba 1.4B being an outlier in terms of poor performance.
+With the exception of TinyLlama at 1-shot performance and Mamba-2 2.7B for 3- and 5-shot performance, the majority of other pretrained models cluster together.
+Instruction tuning greatly smooths ICL performance across both parameter classes.
+While instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all parameters of 270 million and greater show positive ICL abilities.
+The instruction tuned OpenELM 450M model displays particularly impressive ICL abilities, but it is difficult to determine whether it is strictly due to architecture and/or pretraining recipe, or partially due to $143\%$ more total pretraining tokens than Mamba/Mamba-2 and Pythia models.
+For instruction tuned models of greater than 450 million parameters, all SSMs and Transformers show positive ICL abilities, with Mamba-2 2.7B greatly outperforming all other models (both SSM and Transformer) in this class.
+Thus, in terms of ICL as a function of SSM model size, while no clear trend presents itself for pretrained models, ICL appears to emerge for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater.
+6 Discussion and Future Directions
+Using dynamical systems theory, we’ve shown that the recurrent dynamics of Mamba SSMs are robust to small input perturbations.
+We’ve extensively confirmed this result, showing that: a) Mamba inference differences due to mixed-precision align with Transformers, (b) Mamba fine-tuning is significantly more robust to changes due to mixed-precision and PEFT than Transformers, and (c) combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models.
+Using both MPFT and PEFT, we’ve shown that instruction tuning Mamba and Mamba-2 SSMs greatly narrows the pretraining ICL gap on natural language tasks relative to comparable Transformer LLMs.
+In particular, this allows Mamba-2 SSMs to greatly outperform the ICL abilities of a large number of instruction tuned, attention-based LLMs.
+Furthermore, complimentary to recent studies, we’ve shown that ICL for natural language tasks can be characterized as an emergent ability of Mamba and Mamba-2 models of 370 millions parameters or greater.
+There are several avenues for future work.
+In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups and testing the limits of the derived stability results.
+Future studies could further explore what known emergent abilities Mamba SSMs are (and aren’t) capable of, as well as determining new Mamba-specific emergent abilities that are potentially not possible for Transformers to achieve.
+Furthermore, our theoretical contributions open the door for follow up studies, both in terms of extending our stability results to more general error (and adversarial) robustness results, as well as deriving new Mamba SSM-specific LoRA schemes which leverage the implicit weight tying achieved by targeting MambaBlock modules for either PEFT accuracy or efficiency improvements.
+7 Acknowledgments
 We thank Leidos for funding this research through the Office of Technology. Approved for public release 24-LEIDOS-0515-27826.
-Appendix A Mamba stable dynamics proof
-Recall the state-space parameters and equations for the MambaBlock; $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,n\}=[n]$.
-Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{n}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{n}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{n}\in\mathbb{R}{}^{d}$:
+Appendix A Proof of weight-tying using LoRA in the MambaBlock
+Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer.
+E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these individual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer).
+However, we note that the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
+Theorem 1.
+Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 4.
+Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
+Proof.
+Let $r$ be the specified LoRA dimension. Targeting this matrix for LoRA results in the adapter
+$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
+$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
+where $\mathbf{U}\in\mathbb{R}{}^{n\times r}$, $\mathbf{V}\in\mathbb{R}{}^{r\times 3d}$, and $\mathbf{W}$ is frozen during fine-tuning. Thus, for index $[i,j]$,
+$\mathbf{W}^{\prime}[i,j]=$ $\sum_{k=0}^{r-1}\mathbf{U}[i,k]\mathbf{V}[k,j].$
+Recall the form of $\mathbf{W}$:
+$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$
+where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. For index $[t-1,j]$, we thus have
+$\tilde{\mathbf{W}}[t-1,j]=$ $\mathbf{W}[t-1,j]+\mathbf{W}^{\prime}[t-1,j]$
+$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
+Thus, the weights $\mathbf{U}[t-1,:]$ are tied for any parameter $\tilde{\mathbf{W}}[t-1,j],j\in\{1,\dots,3d\}$, which are used to adapt parameters $\bm{\Delta}_{1},\mathbf{B}_{t},$ and $\mathbf{C}_{t}$.
+Appendix B Mamba stable dynamics proof
+Recall the state-space parameters and equations for the MambaBlock; $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}=[T]$.
+Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{T}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{T}\in\mathbb{R}{}^{d}$:
 $\bm{x}_{t}$ $=\bar{\mathbf{A}}_{t}\bm{x}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{u}_{t}$ (5)
@@
 ∎
-Appendix B Proof of weight-tying using LoRA in the MambaBlock
+Appendix C Mamba stable outputs proof
 Theorem 3.
-Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 3.
-Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
+Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time.
+Then small changes to the output $\mathbf{y}_{t}$ are also exponentially non-increasing over discrete time.
 Proof.
-Let $r$ be the specified LoRA dimension. Targeting this matrix for LoRA results in the adapter
-$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
-$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
-where $\mathbf{U}\in\mathbb{R}{}^{n\times r}$, $\mathbf{V}\in\mathbb{R}{}^{r\times 3d}$, and $\mathbf{W}$ is frozen during fine-tuning. Thus, for index $[i,j]$,
-$\mathbf{W}^{\prime}[i,j]=$ $\sum_{k=0}^{r-1}\mathbf{U}[i,k]\mathbf{V}[k,j].$
-Recall the form of $\mathbf{W}$:
-$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$
-where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. For index $[t-1,j]$, we thus have
-$\tilde{\mathbf{W}}[t-1,j]=$ $\mathbf{W}[t-1,j]+\mathbf{W}^{\prime}[t-1,j]$
-$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
-Thus, the weights $\mathbf{U}[t-1,:]$ are tied for any parameter $\tilde{\mathbf{W}}[t-1,j],j\in\{1,\dots,3d\}$, which are used to adapt parameters $\bm{\Delta}_{1},\mathbf{B}_{t},$ and $\mathbf{C}_{t}$.
-Appendix C Experimental Details
+Recall that $\bm{x}_{T}=F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})$. Furthermore, recall from Equations 1 and  2, $\mathbf{y}_{t}=\mathbf{C}_{t}\bm{x}_{t}$, where $\mathbf{C}_{t}$ is diagonal.
+Let
+$\mathbf{y}_{T}$ $=G_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})=\mathbf{C}_{T}\bm{x}_{T}=\mathbf{C}_{T}F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1}).$
+Consider $\varepsilon$-close inputs $(\bm{x}_{t-1},\mathbf{u}_{t})$ and $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$, and their respective outputs $\mathbf{y}_{t}$ and $\mathbf{y}_{t}^{\prime}$.
+Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time.
+That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
+We thus have
+$\max|\mathbf{y}_{t}-\mathbf{y}_{t}^{\prime}|$ $=\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$
+$=\max|\mathbf{C}_{N}F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-\mathbf{C}_{N}F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$
+$\propto\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|,$
+where proportionality follows due to the diagonality of $\mathbf{C}_{N}$ and the vector-absolute value. Thus,
+$\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$ $\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$
+Appendix D Experimental Details
 All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16], version 0.4.2.
@@
 All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0.
+All Mamba-2 models were run using mamba-ssm v2.2.2 using Huggingface checkpoints, e.g., state-spaces/mamba-130m for the smallest model.
 For MPFT, Flash Attention 2.0 [9] via flash_attn 2.5.7 was used for Pythia models.
@@
 For OLMo results, the 336B-token checkpoint was used by specifying revision=step80000-tokens336B.
-Outside of the OpenHermes hyperparameter search, all Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps.
+All Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps.
 Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively.
@@
 Pythia targeted SLL LoRA layers were {dense, embed_in, query_key_value, dense_h_to_4h,dense_4h_to_h}, chosen to balance performance across model sizes.
-All experiments in Tables 1 and  2, Figures 1 and  2(b) were run using a single-GPU Nvidia A10G (24 GB total memory).
-For Pythia and Mamba ALL LoRA experiments in Figure 3, all experiments were run on an A10G, except for Mamba 2.8B, which exceeded A10G memory capacity and was run on an Nvidia H100 (80 GB total memory).
+All experiments were run using a single-GPU Nvidia A10G (24 GB total memory).
+For Pythia, Mamba, and Mamba-2 ALL LoRA experiments in Figure 3, all models followed the same training and PEFT recipes, save for Mamba-2 2.7B which required a LoRA $r$ dimension of 64 to fit in A10G memory.
 Table 2: Learning rate and LoRA dimension $r$ values
-Mamba size Pythia size learning rate LoRA $r$
-130M 160M 1.0e-5 8
-370M 410M 5.0e-5 16
-790M 1B 1.0e-6 32
-1.4B 1.4B 5.0e-6 64
-2.8B 2.8B 5.0e-7 128
-For the hyperparameter search results in Figure 4, all experiments were run using 8 H100 GPUs. SLL LoRA was used for Mamba and Pythia models. The range of hyperparameter values was as follows:
-  • learning rate $\in\{1e-7,2e-7,5e-7,1e-6,2e-6,5e-6,1e-5,2e-5,5e-5,1e-4,2e-4,5e-4,1e-3,2e-3,5e-3\}$
-  • LoRA dimension $r\in\{16,32,64,128,256\}$
-  • warmup steps $\in\{0,1000,2000\}$
-All other hyperparameters followed previous experiments.
+Mamba size Mamba-2 size Pythia size learning rate Mamba/Pythia LoRA $r$ Mamba-2 LoRA $r$
+130M 130M 160M 1.0e-5 8 8
+370M 370M 410M 5.0e-5 16 16
+790M 780M 1B 1.0e-6 32 32
+1.4B 1.3B 1.4B 5.0e-6 64 64
+2.8B 2.7B 2.8B 5.0e-7 128 64
 The Alpaca dataset is freely available for download at https://huggingface.co/datasets/tatsu-lab/alpaca under open-source license CC-by-NC 4.0.
 The OpenHermes dataset is freely available for download at https://huggingface.co/datasets/teknium/OpenHermes-2.5 under open-source license MIT, Apache 2.0, CC.
+Appendix E Mamba-2 mixed-precision inference pretrained LLM performance tables
+Table 3: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference.
+Divergence is averaged over {0, 1, 3, 5}-shot performance.
+Pretrained checkpoints are used for Mamba (M), Mamba-2 (M2), Pythia (P), OLMo [19] (O), and Phi-1.5 [33] (Phi) models.
+Model M M2 P M M2 P M M2 P O M M2 P Phi M M2 P
+Size 130m 130m 160m 370m 370m 410m   790m 780m 1b 1.4b 1.3b 1.4b 1.5b 2.8b 2.7b 2.8b
+FP16 $\mu$ 0.03 0.07 0.35 0.05 0.05 0.06 0.21 0.12 0.05 0.04 0.04 0.04 0.07 0.03 0.15 0.26 0.12
+BF16 $\mu$ 0.05 0.67 1.45 0.20 0.52 0.20 0.66 0.29 0.16 0.13 0.31 0.40 0.13 1.05 1.17 1.02 0.11
+Appendix F Adapted PEFT-layer effects on Mamba-2 zero-shot performance
+As described in Section 3.2, both Mamba and Mamba-2 store the majority of time-varying SSM parameters in a large memory buffer (instantiated as a PyTorch linear layer’s weight matrix), denoted as $\mathbf{W}$.
+For PEFT via LoRA, we ablate the impact of targeting $\mathbf{W}$, as well as the impact of Theorem 3, where time-varying variables accessed through $\mathbf{W}$ were shown to be regularized through weight tying.
+Firstly, we instruction tune Mamba-2 models using LoRA by both adapting $\mathbf{W}$ and adapting all linear layers other than $\mathbf{W}$.
+Mamba-2 MambaBlocks contain two linear layers, thus the former targets only the $\mathbf{W}$ linear layer whereas the latter targets the other linear layer in each block.
+As shown in Table 4, PEFT with only $\mathbf{W}$ targeted within the MambaBlock near uniformly results in better performance than only targeting other layers, with the former outperforming the latter on 32 out of 35 natural language tasks.
+This makes intuitive sense since $\mathbf{W}$ represents the majority of Mamba’s time-varying parameters (as previously mentioned).
+Thus, while previous works [21, 11] have displayed the importance of Mamba’s time-varying parameters for pretraining performance, this verifies the importance of Mamba’s time-varying parameters for instruction tuning performance.
+Table 4: Zero-shot performance for instruction tuned Mamba-2 models where: denotes the large memory buffer $\mathbf{W}$ containing the majority of temporal variables (described in Section 3.2) is targeted for LoRA adaptation, and denotes $\mathbf{W}$ is not adapted.
+The top-performance for each task per model is highlighted in bold .
+Model $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
+ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
+Mamba-2 130M 15.37 45.16 35.41 65.02 47.94 24.91 52.17
+16.95 43.30 35.24 64.85 47.47 24.06 52.17
+Mamba-2 370M 8.03 54.16 46.87 69.53 53.49 27.73 57.14
+8.50 53.74 46.59 70.67 54.80 26.79 55.64
+Mamba-2 780M 5.79 61.61 55.14 71.98 61.11 29.18 60.09
+5.86 61.63 54.94 72.03 60.98 28.41 60.22
+Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
+5.05 65.44 59.84 73.39 64.10 33.11 61.09
+Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
+4.10 69.61 66.60 76.39 69.53 36.26 63.93
+Next, we present a counterintuitive experiment to demonstrate the impact of Theorem 3.
+As before, we instruction tune Mamba-2 models using LoRA, focusing on the two largest model sizes.
+This time, however, we target all linear layers within Mamba-2’s MambaBlock and contrast this with targeting only $\mathbf{W}$ within the MambaBlock.
+Displayed in Table 5, we can see that adapting more SSM parameters does not necessarily lead to improved performance across the board.
+Rather, adapting only $\mathbf{W}$ for Mamba-2 outperforms adapting all linear layers on eight of the 14 natural language tasks.
+The results in Table 5 thus present two cases: (a) regularization: only adapting parameters which lead to weight tying (by Theorem 3) and (b) increased learning capacity: adapting more parameters, at the cost of learning unregularized parameters.
+Through regularization via weight tying [42], the former leads to good generalization.
+In contrast, while adapting more parameters, (b) includes unregularized variables, leading to positive improvements on only a minority of tasks when compared to the fine-tuning of fewer, but regularized, parameters.
+We note that adapting all linear layers is often performed when fine-tuning attention-based LLMs using LoRA [12].
+Thus, a similar regularization-vs-capacity tradeoff for Transformers LLMs may present itself by carefully studying LoRA’s affects when targeting specific linear layers in attention-based architectures.
+Table 5: Zero-shot performance for instruction tuned Mamba-2 models.
+denotes only $\mathbf{W}$ was targeted for LoRA adaptation within the MambaBlock, and thus denotes both linear layers within each MambaBlock were adapted.
+The top-performance for each task per model is highlighted in bold.
+Model Only $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
+ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
+Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
+4.94 62.95 60.06 73.78 64.02 34.81 61.33
+Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
+3.77 68.39 67.51 75.95 69.53 38.31 65.35

### v3

@@
 Abstract
-Mamba [21, 11] state-space models (SSMs) were recently shown to outperform state-of-the-art (SOTA) Transformer large language models (LLMs) across various tasks.
-Despite subsequent widespread adaptation, little work has focused on Mamba LLMs’ amenability for fine-tuning frameworks ubiquitously used for Transformer-based LLMs, e.g., mixed-precision fine-tuning (MPFT) and parameter-efficient fine-tuning (PEFT).
-For the former, it currently remains an open question whether Mamba’s recurrent dynamics are robust to small input changes, such as those encountered during MPFT.
-Using dynamical systems theory (in particular, Lyapunov exponents), we answer this question in the affirmative.
-We empirically validate this result through several experiments, showing that Mamba SSMs are significantly more stable to changes introduced by mixed-precision than comparable Transformers, even when both MPFT and PEFT are combined.
-For PEFT, we show how targeting specific memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus providing both parameter efficient learning and computational savings.
-Finally, with both MPFT and PEFT enabled, we explore the impact of instruction tuning Mamba SSMs for in-context learning (ICL) on natural language tasks.
-While pretrained Mamba and Mamba-2 models only achieve 38% and 82% (respectively) of the ICL improvements of comparable Transformer-based LLMs, we show that instruction tuning allows Mamba models to narrow this gap to 81% and Mamba-2 models to skyrocket over this gap to 132%.
-**footnotetext: Correspondence to john.t.halloran@leidos.com.
+Mamba state-space models (SSMs) have recently outperformed state-of-the-art (SOTA) Transformer large language models (LLMs) in various tasks and been widely adapted.
+However, a major concern for stable learning in recurrent-based deep models (such as SSMs) is the sensitivy of their recurrent dynamics.
+Despite widespread adaptation, the sensitivity of Mamba’s recurrent dynamics under common fine-tuning methods–e.g., mixed-precision fine-tuning (MPFT) and parameter-efficient fine-tuning (PEFT)–remains unexplored.
+Empirically, we show that Mamba LLMs are extremely stable to changes introduced by combinations of MPFT and PEFT, in stark contrast to Transformer LLMs, which we demonstrate may drastically diverge from their respective full-precision counterparts under different combinations of MPFT and PEFT (despite the near-ubiquitous adaptation of these fine-tuning frameworks for attention-based models).
+The demonstrated robustness of Mamba LLMs are due to their recurrent dynamics, which we prove are guaranteed to be stable using dynamical systems theory (in particular, Lyapunov stability).
+We conclude by using MPFT and PEFT to novelly study Mamba LLMs’ in-context learning (ICL) abilities on natural language tasks, thus supplementing other recent work.
 1 Introduction
-Innovating on previous state-space models (SSMs) [22, 10], Mamba [21, 11] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs).
-Upon their introduction in [21], Mamba SSMs were shown to greatly outperform comparable attention-based LLMs [4] across a large number of standard natural language benchmarks.
-Subsequently, pretrained Mamba models have been widely adapted across a large number of data modalities [36, 31, 43, 32], tasks [57, 53], and architectures  [1, 40, 34].
-Despite such widespread adaptation and subsequent research threads [11, 40, 54], little work has been done to understand the amenability of Mamba SSMs for widely used fine-tuning frameworks, such as mixed-precision fine-tuning (MPFT) [38] and parameter-efficient fine-tuning (PEFT) [23, 25].
-MPFT and PEFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [51], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12].
-However, direct application of MPFT for Mamba SSMs is made difficult due to potential sensitivities of Mamba’s state-space dynamics, a common concern for recurrent-based deep models [41].
-To combat this, both [26] and [20] suggest full precision (FP32) may be required to perform stable training for Mamba models.
-Thus, it is currently an open question whether Mamba’s recurrent dynamics are stable in the presence of small input deviations, such as those introduced in MPFT.
-To answer this question, we leverage theory from dynamical systems.
-Deriving and bounding the Lyapunov exponents for both Mamba and Mamba-2 models, we show that small input changes within the SSM layer of either model do not lead to exponentially deviating outputs.
-Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference are on par with those demonstrated by Transformer LLMs, while deviations due to MPFT are significantly more stable than those of comparable Transformers (Section 5).
-Furthermore, this trend continues when MPFT and PEFT are combined, where Mamba SSMs again produce significantly smaller deviations compared to comparable Transformer LLMs.
-For PEFT, we show that by targeting the large memory buffers exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying.
-We show that this leads to extremely efficient PEFT, resulting in up to 2.15 times faster training and 65.5% reduced memory compared to the largest evaluated Mamba model without MPFT or PEFT.
-Furthermore, this allows even the largest (2.8 billion parameter) Mamba LLMs to be fine-tuned on a single GPU with as little as 24GB of onboard memory.
-Finally, using both MPFT and PEFT, we complement existing studies [40, 30] by exploring the ICL capabilities of instruction-tuned Mamba and Mamba-2 models on natural language tasks.
-In particular, the ICL capabilities of both pretrained Mamba and Mamba-2 models lag behind those of comparable Transformer models from the Pythia suite [4]; averaged across five standard natural language benchmarks and foundation model sizes, Mamba and Mamba-2 models only achieve 38% and 82%, respectively, of the performance improvements (relative to zero-shot) of Pythia models.
-However, after instruction-tuning, Mamba models are able to achieve as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers, while Mamba-2 models push this to 132% of the ICL improvements achieved by Pythia models.
-We note that, similar to Transformer foundation models [55], (post) instruction tuning ICL appears as an emergent abilities for Mamba and Mamba-2 SSMs, manifesting for models of size 370 million parameters and greater, while failing to manifest for Mamba and Mamba-2 models of fewer parameter counts.
-We note that herein, when describing a particular foundation model or result, we use the term “Mamba model” to refer to one of the original models released in [21] and “Mamba-2 model” to refer to models released in [11].
-While there are subtle architectural differences between these two SSMs, they share important similarities which allow our theoretical results to extend to both sets of models.
-In particular, Mamba and Mamba-2 models share the same state-space equations, support for SSM matrices, and design scheme of storing the majority of SSM parameters in a large memory buffer.
-Thus, we synonymously use the term MambaBlock to refer to the SSM layer of both Mamba and Mamba-2 models.
+Mamba(Gu & Dao, 2024a) state-space models (SSMs) have been shown to greatly outperform comparable attention-based LLMs (Biderman et al., 2023) across a large number of standard natural language benchmarks.
+Subsequently, pretrained Mamba models have been widely adapted across a large number of data modalities (Liu et al., 2024; Li & Chen, 2024; Quan & Li, 2024; Li et al., 2024), tasks (Xie et al., 2024; Wang et al., 2024a), and architectures  (Anthony et al., 2024; Park et al., 2024; Lieber et al., 2024).
+However, despite such widespread adaptation and subsequent research threads (Dao & Gu, 2024; Park et al., 2024; Wang et al., 2024b), no existing work has sought to understand the stability of learning Mamba SSMs in common LLM fine-tuning frameworks, such as mixed-precision fine-tuning (MPFT).
+For recurrent-based deep models, such as Mamba SSMs, the sensitivity of recurrent dynamics to small changes (e.g., those introduced by mixed-precision) is a primary concern for stable learning (Bengio et al., 1994; Pascanu et al., 2013).
+To determine the stability of time-varying Mamba models, we leverage theory from dynamical systems.
+Under the framework of Lyapunov stability, we show that small input changes within the SSM layer of both Mamba and the more recent Mamba-2 (Dao & Gu, 2024) models do not lead to exponentially deviating outputs.
+Furthermore, we extend these theoretical results to account for deviations in input and latent Mamba states produced by combinations of MPFT and PEFT.
+We empirically validate these theoretical results; for a large number of randomly generated SSM layers, we show that manually adjusting initial latent and input states produces maximum deviations in the output states which exponentially decrease over discrete time.
+Furthermore, by expanding previous divergence performance metrics (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023; Dettmers et al., 2024) and evaluating combinations of MPFT and PEFT, we show that fine-tuned Mamba LLMs do not substantially deviate in performance compared to full-precision full fine-tuning.
+This is demonstrated in stark contrast to comparable Transformer-based LLMs, which we show may drastically differ from their full-precision full fine-tuning counterparts, exhibiting large deviation spikes.
+Over two widely used fine-tuning datasets and two widely used natural language benchmarks, attention-based LLMs Pythia (Biderman et al., 2023) and OpenELM (Mehta et al., 2024) exhibit 9 and 4 large deviation spikes, respectively, whereas Mamba LLMs exhibit zero.
+Thus, Mamba LLMs are significantly more stable to performance deviations under MPFT and PEFT, whereas Transformer LLMs are susceptible to large deviation spikes.
+Lastly, we use stable MPFT and PEFT to explore the ICL capabilities of instruction tuned Mamba and Mamba-2 models on natural language tasks, thus complementing recent studies which evaluated Mamba ICL on synthetic tasks (Park et al., 2024; Lee et al., 2024).
+While the ICL capabilities of pretrained Mamba models lag behind those of comparable Transformer models–Mamba and Mamba-2 models only achieve 38% and 82%, respectively, of the performance improvements (relative to zero-shot) of Pythia models–instruction tuned Mamba and Mamba-2 models improve to 81.5% and 132% of the ICL improvements achieved by Pythia.
+Thus, while pretrained Mamba models’ ICL capabilities lag behind comparable Transformers, we show that instruction-tuning using MPFT and PEFT greatly narrows this gap.
+Our major contributions are summarized as follows:
+  • We derive bounds on the Lyapunov exponents of both Mamba and Mamba-2 models’ SSM equations.
+Using these bounds, we theoretically show that small input changes within the MambaBlock do not lead to exponentially deviating outputs.
+  • Building on the aforementioned bounds, we theoretically show that changes to MambaBlock latent and input states stemming from MPFT and PEFT do not lead to exponentially deviating outputs.
+  • We empirically demonstrate the above theoretical results:
+    • Directly introducing $\varepsilon$ changes to the inputs states of 100 random MambaBlock, we confirm that the resulting output states do not deviate from the unperturbed outputs over discrete time (the maximum deviation exponentially decreases).
+    • For MPFT and PEFT, we measure the performance deviation between FP32 full fine-tuned and MPFT/LoRA Mamba models and compare to attention-based models.
+Across two fine-tuning datasets, two widely used natural language benchmarks, several model sizes, and a large number of MPFT/PEFT configurations, we show that training Mamba LLMs is significantly more stable than comparable Transformer LLMs.
+  • We expand performance divergence measures previously used to study MPFT and PEFT for Transformer LLMs (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023; Dettmers et al., 2024).
+In addition to showing Mamba models are drastically more stable than comparable attention-based LLMs, the expanded divergence metric also demonstrates that Transformer LLMs are susceptible to large deviation spikes.
+  • Using stable MPFT and PEFT, we complement recent studies by examining the ICL capabilities of Mamba/Mamba-2 models evaluated on natural language tasks.
+We show that instruction tuning allows SSMs to perform ICL competitively with comparable Transformer LLMs on natural language tasks.
+Terminology.
+When describing a particular foundation model or result, we use the term “Mamba model” to refer to one of the original models released in Gu & Dao (2024a) and “Mamba-2 model” to refer to models released in Dao & Gu (2024).
+Despite subtle architectural differences, these two SSMs share the same state-space equations and design scheme of storing the majority of SSM parameters in a large memory buffer.
+We thus synonymously use the term MambaBlock to refer to the SSM layer of both Mamba and Mamba-2 models.
 2 Background
@@
 In the current era of extremely large foundation models, both MPFT and PEFT have become ubiquitous tools for the rapid adaptation of Transformer-based LLMs towards specific applications.
-PEFT using adapters [23] allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters.
-Arguably the most widely used such PEFT method is LoRA [25], which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
-To further decrease the computational demands necessary for LLM fine-tuning and inference, MPFT via mixed-precision (i.e., FP16 or BF16[27, 38] and quantized low-precision [12] have proven effective strategies to reduce GPU memory and runtime requirements without deleterious effects on downstream performance [12, 56].
-Additionally, mixed-precision approaches have paved the way for hardware-aware optimizations within the self-attention module [9], greatly mitigating the quadratic complexity of Transformer LLMs.
+PEFT using adapters (He et al., 2021) allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters.
+Arguably the most widely used such PEFT method is LoRA (Hu et al., 2021), which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
+To further decrease the computational demands necessary for LLM fine-tuning and inference, MPFT via mixed-precision (i.e., FP16 or BF16(Kalamkar et al., 2019; Micikevicius et al., 2018) and quantized low-precision (Dettmers et al., 2024) have proven effective strategies to reduce GPU memory and runtime requirements without deleterious effects on downstream performance (Dettmers et al., 2024; Wu et al., 2020).
+Additionally, mixed-precision approaches have paved the way for hardware-aware optimizations within the self-attention module (Dao et al., 2022), greatly mitigating the quadratic complexity of Transformer LLMs.
 Together, PEFT and MPFT have created a rich ecosystem with which varying combinations of these approaches may be used to meet the computational constraints of a given training system.
-We note that post-fine-tuning quantization approaches [13] may be further used to decrease Transformer LLM computational demands, but such approaches are not considered in this work.
 State-space Models.
-Structured state-space sequence (S4) models [22, 14] are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency.
+Structured state-space sequence (S4) models (Gu et al., 2022; Fu et al., 2023) are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency.
 Within the S4 layer, an input signal is discretized and LTI parameters representing the input’s latent dynamics are learned.
 Owing to the S4 block’s latent dynamics being LTI, the S4 block’s output may be thus compactly represented as a single convolution between the input and an SSM convolution kernel (a matrix whose entries are products of LTI learnable parameters resulting from unrolling the state-space equations).
-However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures [21].
+However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures (Gu & Dao, 2024a).
 Innovating on these previous S4 approaches, Mamba utilizes time-varying parameters to model latent dynamics, thus broadening the ability to capture nuanced changes evolving in discrete-time.
-Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations [14].
-To enable hardware efficiency with time-varying SSM parameters, [21] thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states.
-Subsequently, [11] considered the unrolled state-space equations and leveraged tensor contractions (i.e., einsum notation [45]) to efficiently calculate Mamba variables.
-The resulting Mamba-2 foundation models contained significantly larger latent-variable dimensions than the Mamba models of [21], while maintaining efficiency on modern GPU accelerators.
+Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations (Fu et al., 2023).
+To enable hardware efficiency with time-varying SSM parameters, (Gu & Dao, 2024a) thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states.
+Subsequently, Dao & Gu (2024) considered the unrolled state-space equations and leveraged tensor contractions (i.e., einsum notation (Rogozhnikov, 2022)) to efficiently calculate Mamba variables.
+The resulting Mamba-2 foundation models contained significantly larger latent-variable dimensions than the Mamba models of (Gu & Dao, 2024a), while maintaining efficiency on modern GPU accelerators.
 In-context learning.
@@
 Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest.
-Given sufficiently large models and pretraining data [7, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting.
-However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [7, 35].
+Given sufficiently large models and pretraining data (Brown et al., 2020; Wei et al., 2022), Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting.
+However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results (Brown et al., 2020; Liu et al., 2022).
 3 Mamba state-space models
@@
 $\bm{x}_{t}$ $=\bar{\mathbf{A}}_{t}\bm{x}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{u}_{t}$ (1)
-$\mathbf{y}_{t}$ $=\bar{\mathbf{C}}_{t}\bm{x}_{t},$ (2)
+$\mathbf{y}_{t}$ $=\mathbf{C}_{t}\bm{x}_{t},$ (2)
 where $\bar{\bm{\Delta}}_{t}=\texttt{softplus}(\texttt{Linear}(\bm{\Delta}_{t}))\in\mathbb{R}{}^{d\times d}$, $\bar{\mathbf{A}}_{t}=\exp{(\bar{\bm{\Delta}}_{t}\mathbf{A})}$ and $\bar{\mathbf{B}}_{t}=\mathbf{A}^{-1}(\bar{\mathbf{A}}-\mathbf{I})\mathbf{B}_{t}$.
 In practice, $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are diagonal matrices.
+Due to the time-variance of Equations 1 and  2, previous SSM stability results are no longer applicable (discussed further in Appendix E).
+The Mamba foundation models were pretrained in full FP32 precision and, subsequently, official implementations have cautioned against training in reduced precision (Gu & Dao, 2024b; Huggingface, 2024).
+Thus, how sensitive the MambaBlock’s recurrent dynamics are remains an open question.
+We next answer the latter using theory from dynamical systems.
 3.1 Stable dynamics in the MambaBlock
-The Mamba foundation models were pretrained in full FP32 precision.
-Consequently, official Mamba implementations have cautioned against fine-tuning or training in reduced precision [20, 26], with potential sensitivities of MambaBlock recurrent dynamics remaining an open question.
-We answer the latter using theory from dynamical systems.
-For Mamba’s discrete dynamic system in Equations 1 and  2, define
+For Mamba’s discrete dynamical system in Equations 1 and  2, define
 $\bm{x}_{t}$ $=F_{\theta}(\bm{x}_{t-1},\mathbf{u}_{t}),$ (3)
@@
 $\bm{x}_{T}$ $=F_{\theta}(F_{\theta}(\dots F_{\theta}(\bm{x}_{0},\mathbf{u}_{1})))\coloneq F_{\theta}^{T-1}(\bm{x}_{0},\mathbf{u}_{1}).$
-The rate of divergence between two scalar $\varepsilon$-close inputs to a discrete dynamical system is bounded by the system’s maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ [39].
-Given $\lambda_{\mathtt{max}}{}$ and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$, the maximum deviation between these points grows as [28, 47]:
-$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\lambda_{\mathtt{max}}{})}).$
+The rate of divergence between two scalar $\varepsilon$-close inputs to a discrete dynamical system is bounded by the system’s maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ (Mikhaeil et al., 2022).
+Given $\lambda_{\mathtt{max}}{}$ and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$, the maximum deviation between these points grows as (Laffargue et al., 2013; Sayama, 2015):
+$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\lambda_{\mathtt{max}}{})}).$
 Thus, when $\lambda_{\texttt{max}}>0$, nearby trajectories exponentially separate and, when $\lambda_{\texttt{max}}\leq 0$, nearby trajectories ultimately converge to the same fixed point or periodic cycles.
@@
 That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
-The proof of Theorem 1 is available in Appendix B, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
+The proof of Theorem 1 is available in Appendix A, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
 Thus, the latent states of Mamba and Mamba-2 models are stable under small input changes.
@@
 Then small changes to the output $\mathbf{y}_{t}$ are also exponentially non-increasing over discrete time.
-The proof of Theorem 2 is available in Appendix C. Thus, by Theorems 1 and  2, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during recurrency.
-3.1.1 Consequences for automatic mixed-precision
+The proof of Theorem 2 is available in Appendix B. Thus, by Theorems 1 and  2, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during recurrency.
+3.1.1 Divergence bounds for automatic mixed-precision and PEFT
+For an arbitrary MambaBlock, let $\bm{x}_{0},\dots,\bm{x}_{T}$ and $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ be the trajectory of latent and input states under full-precision fine-tuning.
 During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16).
 During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating.
-In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
-For a discrete dynamical system with $\lambda_{\texttt{max}}>0$, changes due to AMP compound after repeated expansion of the recurrent state, thus leading to exponential deviations between quantities calculated using mixed- versus full-precision.
-We note that Transformers are not recurrent, and thus not susceptible to such issues.
-Yet, just as differences introduced by quantization/mixed-precision produce output differences in Transformer results, differences are expected in Mamba results using different precision strategies.
-However, by Theorem 1, such differences do not exponentially compound over discrete-time within the MambaBlock.
-3.2 Hardware-aware optimizations and PEFT
-As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [10] are no longer applicable.
-However, by diagonality, each dimension may be computed in parallel.
-Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [6].
-[21] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated.
-Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{T\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
-$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$ (4)
-where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on.
-The customized Mamba prefix scan kernel heavily relies on this memory layout to optimize the access pattern of $\mathbf{W}$ in Equations 1 and 2.
-Similarly, Mamba-2 stores diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ in a large memory buffer $\mathbf{W}$.
-However, rather than utilizing the underlying recurrence to directly compute the hidden state and output at each time-step, [11] consider the matrix resulting from unrolling Equation 2 across all $t$.
-Mamba-2 models thus leverage the structure of the ensuing semiseparable matrix to calculate $\bm{x}_{1},\dots,\bm{x}_{T}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ using tensor contractions, which are highly optimized on modern hardware accelerators.
-When fine-tuning using LoRA, low-rank matrices are used to adapt the frozen weight matrices of targeted linear layers.
-The importance of $\mathbf{W}$ for both Mamba and Mamba-2 models makes it a primary candidate for LoRA adaptation.
-In such cases, selecting $\mathbf{W}$ for LoRA adaptation results in the following:
+MFPT thus introduces changes to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish) and latent states $\bm{x}_{0},\dots,\bm{x}_{T}$ through $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), up/downcasting of parameter updates, and the aforementioned changes to the input states.
+Potentially in concert with MPFT, PEFT via LoRA induces changes by learning low-rank matrices $\hat{\mathbf{A}}_{t},\hat{\mathbf{B}}_{t},\hat{\mathbf{C}}_{t}$
+$\hat{\bm{x}}_{t}$ $=(\bar{\mathbf{A}}_{t}+\hat{\mathbf{A}}_{t})\bm{x}_{t-1}+(\bar{\mathbf{B}}_{t}+\hat{\mathbf{B}}_{t})\mathbf{u}_{t}$
+$\hat{\mathbf{y}}_{t}$ $=(\mathbf{C}_{t}+\hat{\mathbf{C}}_{t})\bm{x}_{t}$
+Let $\bm{x}_{0}^{\prime},\dots,\bm{x}_{T}^{\prime}$ and $\mathbf{u}_{1}^{\prime},\dots,\mathbf{u}_{T}^{\prime}$ be the trajectory of latent and input states under MPFT and/or PEFT.
+We thus have the following two theorems:
 Theorem 3.
-Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 4.
-Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
-The proof of Theorem 3 is available in Appendix A. The specific affects of both targeting $\mathbf{W}$ for fine-tuning and Theorem 3’s impact on generalization are ablated in Appendix F.
-4 Related Work
-Recent work has sought to understand how to efficiently increase Mamba’s hidden-state dimension by restructuring SSM operations using tensor contractions [11], leading to Mamba-2 models.
-A separate line of work has sought to understand the in-context learning (ICL) capabilities of Mamba LLMs when trained from scratch for specific tasks [40, 30].
-Another line of recent work has sought to understand how hybrid Mamba-Transformer models may be directly distilled from Transformer models [54].
-However, to the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or sought to understand the effects of LoRA adaptation on modules within the MambaBlock.
-Lyapunov exponents have previously been considered for classic RNN structures (e.g., vanilla RNNs, LSTMs, GRUs, PLRNNs, etc.) [39, 52], to determine when such models exhibit chaotic dynamics and the impact on the exploding/vanishing gradient phenomena***We note that this continues a long line of research exploring RNNs sensitivity to initial conditions and their subsequent ability to produce chaotic output [44, 29, 2, 3], although previous work did not leverage Lyapunov exponents..
-For more recent S4 neural models, [18] used Hurwitz matrices to characterize the numerical stability of linear time-invariant (LTI) S4 models.
-However, such analysis is not applicable to time-varying models, such as Mamba, nor does it characterize the effects of sensitive dependence on initial conditions (e.g., divergence of two $\varepsilon$ close inputs).
-To the best of our knowledge, no previous works have used Lyapunov exponents to explore the effects of mixed-precision on recurrent neural models or Mamba architectures.
-As previously noted, recent works [40, 30] have studied Mamba’s ability to perform ICL by training Mamba models for specific tasks.
-Such tasks include logistic regression, decision trees, and learning other simple function classes, following the work of [17].
-We emphasize that, in this set up, relatively small Mamba models–33 million and 90 million parameters for [30] and [40], respectively–are trained from scratch for every evaluated task.
-Indeed, [40] notes that subsequent work is necessary to understand Mamba’s ICL capabilities for language modeling using standard natural language benchmarks, as well as for larger model sizes.
-Thus, our study of both the pretrained and instruction tuned ICL capabilities of Mamba/Mamba-2 LLMs for natural language tasks are complimentary to previous works.
+Let $\epsilon_{t}^{u}=\max|\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime}|$, i.e., the maximum, positive scalar-difference between the $t$th MPFT/PEFT and full fine-tuning input state.
+Similarly, let $\epsilon_{t}^{x}=\max|\bm{x}_{t}-\bm{x}_{t}^{\prime}|$ and $\epsilon^{*}=\max\bigcup_{t=0}^{T}\{\epsilon_{t}^{u},\epsilon_{t}^{x}\}$ (where we define $\mathbf{u}_{0}=0$).
+We thus have
+$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\zeta)}),$
+for some scalar $\zeta\leq 0$. Thus, differences introduced by AMP and/or PEFT for Mamba models do not exponentially compound over discrete-time.
+Theorem 4.
+Let $\mathbf{y}_{t}^{\prime}$ be the trajectory of output states produced by $\bm{x}_{0}^{\prime},\dots,\bm{x}_{T}^{\prime}$ and $\mathbf{u}_{1}^{\prime},\dots,\mathbf{u}_{T}^{\prime}$ under MPFT and/or PEFT.
+Then we thus have
+$\max|\mathbf{y}_{T}-\mathbf{y}_{T}^{\prime}|\in\mathcal{O}(\epsilon^{*}\exp{(T\zeta)}),$
+for some scalar $\zeta\leq 0$. Thus, small changes to the output $\mathbf{y}_{t}$ resulting from MPFT and/or PEFT are exponentially non-increasing over discrete time.
+The proof of Theorem 3 is available in Appendix C.
+Theorem 4 may be derived by directly applying the proof technique of Theorem 2 to the results of Theorem 3.
+Thus, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during AMP and MFPT.
 4 Experiments
-To demonstrate the implications of Theorem 2, we explore the performance difference between running inference with full-precision pretrained weights and using mixed-precision (FP16 and BF16) weights.
-Model performance is measured as percent accuracy using the MMLU dataset [24].
+4.1 Non-divergent MambaBlock dynamics
+We explore the implications of Theorem 2 for arbitrary MambaBlocks.
+For $n$ random trials and various $\varepsilon$, we randomly initialize the inputs (i.e., $u_{t}$ for $t=1,\dots,T$) and parameters (i.e., $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}$) of a MambaBlock with $T=2048$ and latent-variable dimension $d=64$.
+For each trial and $\varepsilon$, we first calculate the MambaBlock outputs $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ (where $\mathbf{y}_{t}=F_{\theta}^{t}(\bm{x}_{0},\mathbf{u}_{1})$).
+We then perturb the initial states $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)=(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})$ and calculate the subsequent perturbed model outputs $\mathbf{y}_{1}^{\prime},\dots,\mathbf{y}_{T}^{\prime}$ (where $\mathbf{y}_{t}^{\prime}=F_{\theta}^{t}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})=F_{\theta}^{t}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}^{\prime}+\varepsilon)$).
+The maximum deviation is then calculated for each time point, i.e., $\max|y_{t}-y_{t}^{\prime}|=\max|F_{\theta}^{t}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{t}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|$.
+Figure 1 shows both the mean maximum deviation per time averaged across trials, as well as the standard deviation, for $\varepsilon=0.1$.
+As we can see, the maximum divergence starts out proportional to the change in initial conditions, but quickly decays and converges to the original output trajectory.
+Thus, changes to a MambaBlocks inputs do not compound over discrete time.
+Additional experiments for $\varepsilon\in\{0.01,0.001\}$ are available in Appendix D and demonstrate the same dynamics.
+Figure 1: For 100 random MambaBlocks, initial latent and input states $(\bm{x}_{0},\mathbf{u}_{0})$ are deviated by $\varepsilon$.
+The resulting maximum deviation in output states is computed and shown to exponentially decrease over discrete time.
+$\mu$ denotes the maximum absolute deviation at discrete time $t$ averaged over all trials, while $\sigma$ denotes the respective standard deviation at each discrete time.
+4.2 Non-divergent Mamba MPFT and PEFT
+We explore the implications of Theorem 4 for combinations of MPFT and PEFT on pretrained Mamba models.
+We focus on utilizing LoRA (Hu et al., 2021), which is arguably the most widely used PEFT framework for LLMs.
+Using the Alpaca dataset (Taori et al., 2023), Mamba 160M, 410M, and 790M models are fine-tuned for three epochs with a maximum sequence length of 512.
+We denote the targeting of all linear layers (ALL) for LoRA as ALL LoRA, the targeting of a subset of linear layers (SLL) for LoRA as SLL LoRA (i.e., $\bm{\Delta}_{t},\mathbf{B}_{t}$, and $\mathbf{C}_{t}$ parameters), and no adapters as Full (i.e., full fine-tuning).
+Each fine-tuning run occurred on a single Nvidia A10G GPU (24 GB total memory).
+To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size.
+While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules.
+For comparison, two Transformer-based LLM families of similar parameter counts are fine-tuned using the same experimental setup: Pythia (sizes 160M, 410M, and 1B) and OpenELM (Mehta et al., 2024) (sizes 270M and 450M).
+The training recipe for all models was adapted from Tunstall et al.
+(2023), with the AdamW_torch optimizer and a cosine annealing schedule.
+Further experimental details are available in Appendix F.
+(a) Mean full-precision (FP32) divergence in MMLU performance.
+(b) Mean full-precision (FP32) divergence in Winogrande performance.
+Figure 2: Mean full-precision (FP32) divergence in performance for Mamba, Pythia, and OpenELM models.
+Models are fine-tuned over the Alpaca dataset using different combinations of MPFT and PEFT.
+Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
+All models were evaluated using the LM evaluation harness from Eleuther AI (Gao et al., 2023).
+Model performance is measured as percent accuracy using the MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021) datasets.
 The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy.
 Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy.
-Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [4], OLMo [19] 336B-token checkpoint, and Phi 1.5 [33]).
-We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers.
-To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory).
-All models were evaluated using the LM evaluation harness from Eleuther AI [16].
-Further experimental details are available in Appendix D.
-The results are available in Table 1.
-Table 1: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference.
-Divergence is averaged over {0, 1, 3, 5}-shot performance.
-Pretrained checkpoints are used for Mamba (M), Pythia (P), OLMo [19], and Phi-1.5 [33] (Phi) models.
-Model M P M P M P OLMo M P Phi M P
-Size 130M 160M 370M 410M  790M 1B 1.4B 1.5B 2.8B
-FP16 $\mu$ 0.03 0.35 0.05 0.06 0.21 0.05 0.04 0.04 0.07 0.03 0.15 0.12
-BF16 $\mu$ 0.05 1.45 0.20 0.20 0.66 0.16 0.13 0.31 0.13 1.05 1.17 0.11
-From Table 1, inferencing in Pythia using FP16 and BF16 result in an average 0.13 and 0.41 full-precision divergence, respectively.
-Mamba displays similar averages in comparison: inferencing in Mamba using FP16 and BF16 result in an average 0.10 and 0.48 divergence, respectively.
-Interestingly, both SSM and Transformer architectures exhibit large divergence spikes–i.e., mean divergence greater than a percentage point–when using BF16, which occurs once for Mamba and Phi 1.5 models and twice for Pythia models.
-Due to space constraints, Mamba-2 results for the same experiment are included in Appendix Table 3.
-We note that, for comparable model sizes and mixed-precision, Mamba-2 models follow identical divergence trends as respective Mamba models.
-In the following, we show that the observed large divergence spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
-5.1 Mamba SSMs are stable learners
-We next explore the implications of Theorem 1 on fine-tuning, wherein mixed-precision is especially critical; MPFT combined with PEFT adapters have been shown to drastically reduce Transformer fine-tuning times [12].
-We are thus interested in the divergence between Mamba models fully fine-tuned (i.e., no adapters, all model weights are trained) in full-precision and models fine-tuned using mixed-precision and/or PEFT adapters.
-We focus on utilizing LoRA [25], which is arguably the most widely used PEFT framework for LLMs.
-Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba, Pythia, and OpenELM models.
-Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT.
-Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
-Using the Alpaca dataset [48], Mamba 160M, 410M, and 790M models are fine-tuned for three epochs with a maximum sequence length of 512.
-We denote the targeting of all linear layers (ALL) for LoRA as ALL LoRA, the targeting of a subset of linear layers (SLL) for LoRA as SLL LoRA, and no adapters as Full (i.e., full fine-tuning).
-Both ALL and SLL LoRA adapt the large memory buffer described in Theorem 3.
-Each fine-tuning run occurred on a single A10G GPU.
-To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size.
-While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules.
-For comparison, two Transformer-based LLM families of similar parameter counts are fine-tuned using the same experimental setup: Pythia (sizes 160M, 410M, and 1B) and OpenELM [37] (sizes 270M and 450M).
-The training recipe for all models was adapted from [50], with the AdamW_torch optimizer and a cosine annealing schedule.
-Further experimental details are available in Appendix D.
-For each Mamba, Pythia, and OpenELM model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy.
+We thus refer to a divergence greater than unity as a large deviation spike.
+Mean divergence results for all models are displayed in Figure 2(a) and Figure 2(b) for MMLU and Winogrande, respectively.
 Across mixed-precisions and adapter settings, Mamba displays smaller divergences than both Pythia and OpenELM models.
-E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia and 0.54 for OpenELM.
-Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia and 0.33 for OpenELM.
+E.g., for Winogrande evaluations, Alpaca fine-tuning with Mamba models is an average 2.9 and 2.4 times smaller in mean divergence than Pythia and OpenELM models, respectively.
+For MMLU evaluations, Alpaca fine-tuning with Mamba models is an average 2.6 times smaller in mean divergence than both Pythia and OpenELM models.
 Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning, in contrast to both Pythia and OpenELM models.
-(a) Average tokens-per-second $\uparrow$.
-(b) Maximum memory-per-token $\downarrow$.
-Figure 2: Timing and memory usage calculated Mamba model-sizes and PEFT combinations.
-Each model was trained using the Alpaca dataset dataset for three epochs and maximum sequence length 512.
-For each PEFT combination, the batch size was tuned to maximize GPU occupancy.
-Full fine-tuning exceeds available GPU memory (24 GB) for models greater than 790 million parameters.
-Hardware throughput and memory-utilization improvements.
-With stable dynamics and observed divergences smaller than comparable Transformers, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs.
-To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset.
-However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
-For each MPFT and PEFT configuration, the average tokens-per-second (ATPS) is calculated as the total tokens used for fine-tuning divided by total training time, and the maximum memory-per-token (MMPT) is calculated as the maximum GPU memory utilization incurred (over the entire fine-tuning run) divided by the total number of tokens in each mini-batch.
-Results are plotted in Figure 2(b).
-Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 2(b).
-Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%.
-Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs.
-Furthermore, while full fine-tuning is no longer possible on a single A10G for Mamba models greater than 790 million parameters, MPFT and PEFT allow training Mamba models up to 2.8 billion parameters on GPUs with as little as 24 GB onboard memory.
-5.2 Instruction tuning impact on Mamba ICL for natural language tasks
-Using both MPFT and PEFT, we next explore how instruction tuning affects Mamba and Mamba-2 ICL performance on natural language tasks.
-All Mamba and Mamba-2 pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples).
-We use the training recipe of  [50], which includes BF16 utilization.
-Zero and few-shot performance is evaluated using five standard natural language benchmarks: HellaSwag [58], PIQA [5], Arc-E [8], Arc-C [8], and WinoGrande [46].
+We note that the results in Figure 2 display the fine-tuning of 72 LLMs and 576 few-shot evaluations.
+In Appendix H, these experiments are expanded to include the LIMA fine-tuning dataset (Zhou et al., 2024) as well as the standard deviations for all mean deviation results.
+Across all experiments, Mamba models are significantly more stable for MPFT/PEFT compared to Transformer-based LLMs.
+E.g., for MMLU evaluations, LIMA fine-tuning with Mamba models is an average 7 and 3.3 times smaller in mean divergence than Pythia and OpenELM models, respectively.
+For Winogrande evaluations, LIMA fine-tuning with Mamba models is an average 4.1 and 3.3 times smaller in mean divergence than Pythia and OpenELM models, respectively.
+Across the two fine-tuning datasets, two benchmarks, and six evaluated MPFT/PEFT combinations, Pythia and OpenELM produce 9 and 4 large deviation spikes, respectively, while Mamba produces zero.
+4.3 Hyperparameter tuning robustness of Mamba models
+We empirically demonstrate that Mamba models are robust to the choice of PEFT hyperparemters.
+Using the training recipe of  Tunstall et al.
+(2023) as a basis, we conduct an extensive hyperparameter search across the learning rate, LoRA dimension, and number of warmup steps.
+From the Cartesian-product of these three parameters, 150 hyperparameter configurations were sampled and used to fine-tune Mamba 370M over the Openhermes dataset (Teknium, 2024), which consists of 242,000 supervised samples.
+For comparison, Pythia 410M is similarly fine-tuned using the same set of 150 hyperparameter configurations.
+(a) Mamba 370M
+(b) Pythia 410M
+Figure 3: Fine-tuning hyperparameter search for OpenHermes.
+Each point is a different hyperparameter configuration.
+SLL LoRA was used for both models.
+The $x$-axis is the learning rate, the $y$-axis is resulting MMLU 5-shot performance, bubble size is the LoRA dimension, and the color is the number of warmup steps $\in\{0,1\mbox{k},2\mbox{k}\}$.
+The MMLU 5-shot performance for each of the 150 Mamba and Pythia fine-tuned models is displayed in Figure 3.
+Pythia 410M is capable of higher performance than Mamba 370M, where the average accuracy for the former and the latter are 26.5% and 24.8%, respectively.
+However, Mamba 370M is much more robust to the choice of hyperparameters, with a difference of 1.5% between the minimum (23.3%) and maximum (24.8%).
+In contrast, Pythia 410M fine-tuned models display a large performance difference of 4.7% between the minimum (22.9%) and maximum (27.6%).
+4.4 Stable instruction tuning improves Mamba ICL for natural language tasks
+With both stable MPFT and PEFT, we determine the impact of instruction tuning Mamba and Mamba-2 models on ICL for natural language tasks.
+We instruction fine-tuned Mamba and Mamba-2 pretrained models using ALL LoRA, the OpenHermes dataset, and the training recipe described in Section 4 (with BF16).
+Zero and few-shot performance are evaluated using the same five standard natural language benchmarks used in Section 4.3.
 ICL performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS).
@@
 (d) Mamba ALL LoRA
-(e) Mamba2 ALL LoRA
+(e) Mamba-2 ALL LoRA
 (f) Pythia ALL LoRA
 Figure 4: Instruction tuning narrows the ICL gap between Mamba and Pythia, and creates a gap from Pythia to Mamba-2 models.
-ALL LoRA models were instruction tuned on the OpenHermes [49] dataset for one epoch.
+ALL LoRA models were instruction tuned on the OpenHermes (Teknium, 2024) dataset for one epoch.
 Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks.
@@
 However, after instruction tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase.
-Similarly, the majority of Mamba-2 models larger than Mamba 130M greatly improve in ICL performance.
+Similarly, the majority of Mamba-2 models larger than Mamba-2 130M greatly improve in ICL performance.
 Thus, while pretrained Mamba and Mamba-2 models are only capable of 38% and 82% (respectively) of the AIPSS compared to similar pretrained Pythia models, instruction tuned Mamba and Mamba-2 models are capable of 81.5% and 133% of the AIPSS relative to similarly fine-tuned Pythia models.
-We note that a significant difference between Mamba and Mamba-2 models is the larger (by a factor of four) latent dimension.
-5.3 ICL is an emergent ability of instruction tuned Mamba SSMs
-We next study the emergent behavior (as a function of model size) of Mamba/Mamba-2 SSMs’ ICL abilities on natural language tasks by comparing to a larger number of Transformer-based LLMs of varying sizes.
-We compare to OpenELM [37] (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B [59], and OLMo 1.2B [19].
-To limit the emergent effects on both parameter size and pretraining token counts, we did not evaluate models greater than 2.8 billion parameters and chose open-source checkpoints as close as possible to the 300 billion total pretraining tokens used for Mamba, Mamba-2, and Pythia models.
-Thus, pretraining token counts for OpenELM, TinyLlama, and OLMo models were 429 billion, 503 billion, and 336 billion, respectively.
-We note this potentially biases ICL performance in favor of the newly evaluated Transformer-based LLMs, and that direct comparisons between Mamba, Mamba-2, and Pythia are the most fair (as these three classes of models were all pretrained on the same dataset for the same number of total pretraining tokens).
-We repeat the experiments from Figure 3, where we evaluate the pretrained and instruction tuned ICL capabilities of all models.
-To understand the critical role of parameter counts, we group all models into two classes: LLMs containing 450 million parameters or less, and LLMs containing greater than 450 million parameters.
-ICL performance measured by AIPSS is displayed in Figure 4.
-[…]
-[…]
-[…]
-[…]
-Figure 4: Instruction tuning improves Mamba-2 ICL performance past Transformer LLMs.
-ALL LoRA models were instruction fine-tuned on the OpenHermes dataset for one epoch.
-Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks: HellaSwag, PIQA, Arc-E, Arc-C, and WinoGrande.
-From Figure 4, it is clear that pretrained SSMs and Transformers of parameter counts 270 million and less display slight or detrimental ICL abilities (i.e., few-shot performance is worse than zero-shot).
-For models of greater than 450 million parameters, the majority of SSMs and Transformers display positive ICL abilities, with Mamba 1.4B being an outlier in terms of poor performance.
-With the exception of TinyLlama at 1-shot performance and Mamba-2 2.7B for 3- and 5-shot performance, the majority of other pretrained models cluster together.
+In addition to Pythia, we compare instruction tuned Mamba models’ ICL abilities to other state-of-the Transformer LLMs of comparable sizes and pretraining token counts: OpenELM (Mehta et al., 2024) (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B (Zhang et al., 2024), and OLMo 1.2B (Groeneveld et al., 2024).
+We additionally group models by parameter count, to understand ICL as an emergent behavior of fine-tuned Mamba models.
+To further explore the emergence of this ability for Mamba SSMs, we include recent Mamba-specific PEFT methods (Yoshimura et al., 2025).
+Displayed in Appendix G, pretrained SSMs and Transformers of parameter counts 270 million and less display slight or detrimental ICL abilities (i.e., few-shot performance is worse than zero-shot).
+For models of greater than 450 million parameters, the majority of SSMs and Transformers display positive ICL abilities, with the exception of Mamba 1.4B.
 Instruction tuning greatly smooths ICL performance across both parameter classes.
-While instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all parameters of 270 million and greater show positive ICL abilities.
-The instruction tuned OpenELM 450M model displays particularly impressive ICL abilities, but it is difficult to determine whether it is strictly due to architecture and/or pretraining recipe, or partially due to $143\%$ more total pretraining tokens than Mamba/Mamba-2 and Pythia models.
-For instruction tuned models of greater than 450 million parameters, all SSMs and Transformers show positive ICL abilities, with Mamba-2 2.7B greatly outperforming all other models (both SSM and Transformer) in this class.
-Thus, in terms of ICL as a function of SSM model size, while no clear trend presents itself for pretrained models, ICL appears to emerge for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater.
-6 Discussion and Future Directions
-Using dynamical systems theory, we’ve shown that the recurrent dynamics of Mamba SSMs are robust to small input perturbations.
-We’ve extensively confirmed this result, showing that: a) Mamba inference differences due to mixed-precision align with Transformers, (b) Mamba fine-tuning is significantly more robust to changes due to mixed-precision and PEFT than Transformers, and (c) combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models.
-Using both MPFT and PEFT, we’ve shown that instruction tuning Mamba and Mamba-2 SSMs greatly narrows the pretraining ICL gap on natural language tasks relative to comparable Transformer LLMs.
-In particular, this allows Mamba-2 SSMs to greatly outperform the ICL abilities of a large number of instruction tuned, attention-based LLMs.
-Furthermore, complimentary to recent studies, we’ve shown that ICL for natural language tasks can be characterized as an emergent ability of Mamba and Mamba-2 models of 370 millions parameters or greater.
+Thus, while instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all models of 270 million parameters and greater show positive ICL abilities, with Mamba-2 2.7B closely matching (or exceeding) the ICL ability of the top performing Transformer LLM per-shot.
+Furthermore, in terms of an emergent ability (Wei et al., 2022), ICL emerges for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater, while no clear trend presents itself for pretrained models.
+5 Conclusions and Future Directions
+Using dynamical systems theory, we’ve shown that the recurrent dynamics of Mamba SSMs are stable to small input perturbations.
+We’ve extensively confirmed this result, showing that Mamba training is significantly more robust to changes introduced during MPFT and PEFT than Transformer-based LLMs across widely-used fine-tuning datasets and benchmarks.
+Furthermore, by including many more few-shot evaluations in our divergence metric than previous work, our stability studies uncovered that Transformer LLMs are susceptible to large deviation spikes during MPFT alone, as well as when MPFT is combined with PEFT.
+As MPFT and PEFT are two of the most common LLM fine-tuning frameworks, and as lower precision formats (e.g., NF4 and FP8) become more widely used for fine-tuning, we advocate for more extensive stability analysis (e.g., including several few-shot evaluations when calculating divergence) to expose large deviation spikes.
+In order to mitigate large deviation spikes, we’ve shown that, unlike Transformer LLMs, Mamba models are extremely stable and do not exhibit this phenomena.
+Furthermore, in addition to stable MPFT and PEFT, we’ve shown that a core LLM reasoning ability (ICL) of Mamba models significantly improves with instruction tuning; i.e., we’ve shown that while pretrained Mamba models’ ICL capabilities lag behind those of pretrained LLMs on natural language tasks, instruction tuned Mamba models’ rival (and surpass, in many Mamba-2 cases) the ICL capabilities of instruction tuned Transformers.
+Thus, Mamba models are extremely stable learners with the potential to display the ICL capabilities of attention-based models once instruction tuned.
 There are several avenues for future work.
-In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups and testing the limits of the derived stability results.
-Future studies could further explore what known emergent abilities Mamba SSMs are (and aren’t) capable of, as well as determining new Mamba-specific emergent abilities that are potentially not possible for Transformers to achieve.
-Furthermore, our theoretical contributions open the door for follow up studies, both in terms of extending our stability results to more general error (and adversarial) robustness results, as well as deriving new Mamba SSM-specific LoRA schemes which leverage the implicit weight tying achieved by targeting MambaBlock modules for either PEFT accuracy or efficiency improvements.
-7 Acknowledgments
-We thank Leidos for funding this research through the Office of Technology. Approved for public release 24-LEIDOS-0515-27826.
-Appendix A Proof of weight-tying using LoRA in the MambaBlock
-Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer.
-E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these individual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer).
-However, we note that the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
-Theorem 1.
-Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 4.
-Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
-Proof.
-Let $r$ be the specified LoRA dimension. Targeting this matrix for LoRA results in the adapter
-$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
-$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
-where $\mathbf{U}\in\mathbb{R}{}^{n\times r}$, $\mathbf{V}\in\mathbb{R}{}^{r\times 3d}$, and $\mathbf{W}$ is frozen during fine-tuning. Thus, for index $[i,j]$,
-$\mathbf{W}^{\prime}[i,j]=$ $\sum_{k=0}^{r-1}\mathbf{U}[i,k]\mathbf{V}[k,j].$
-Recall the form of $\mathbf{W}$:
-$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$
-where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. For index $[t-1,j]$, we thus have
-$\tilde{\mathbf{W}}[t-1,j]=$ $\mathbf{W}[t-1,j]+\mathbf{W}^{\prime}[t-1,j]$
-$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
-Thus, the weights $\mathbf{U}[t-1,:]$ are tied for any parameter $\tilde{\mathbf{W}}[t-1,j],j\in\{1,\dots,3d\}$, which are used to adapt parameters $\bm{\Delta}_{1},\mathbf{B}_{t},$ and $\mathbf{C}_{t}$.
-Appendix B Mamba stable dynamics proof
+In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods (Dettmers et al., 2024) would further decrease the hardware needed to train Mamba models, while providing additional speedups and testing the limits of the derived stability results.
+Furthermore, our theoretical contributions open the door for several follow up studies.
+E.g., additional results building off Theorems 3 and  4 may tackle generalization error bounds and privacy fine-tuning (by considering a language modeling framework and calculating upper bounds under distributions of $\varepsilon$), and Mamba-specific decoding schemes (where new algorithms may exploit the derived deviation bounds ensure next-token predicitions lie within a deviation tolerance from the true optimal decoding).
+6 Relations to Previous Work
+MPFT divergence for LLMs.
+Previous work has studied the performance difference for attention-based LLMs under low-precision inference (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023) and fine-tuning (Dettmers et al., 2024).
+However, these works only measured divergence using single-shot performance (i.e., zero-shot in  (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023), 5-shot for MMLU and zero-shot otherwise in  (Dettmers et al., 2024)).
+In contrast, we report the difference in model performance as the mean divergence between the full-precision and mixed-precision models, averaged over the accuracy of four different few-shot evaluations.
+Thus, in contrast to previous work, our inclusion of multiple shots: a) uncovers significant divergence in widely used LLMs even at widely used mixed-precisions (FP16 and BF16), and (b) more rigorously tests mixed-precision training’s effect on one of LLMs’ most important reasoning abilities (ICL).
+We are not aware of any previous works which have attempted to understand the performance impact of MPFT on Mamba models.
+Lyapunov stability and RNNs/SSMs.
+Lyapunov exponents have previously been considered (Mikhaeil et al., 2022; Vogt et al., 2022) for classic RNN structures (e.g., vanilla RNNs, LSTMs, GRUs, PLRNNs, etc.), to determine when such models exhibit chaotic dynamics and the impact on the exploding/vanishing gradient phenomena [^1] We note that this continues a long line of research exploring RNNs sensitivity to initial conditions and their subsequent ability to produce chaotic output (Ribeiro et al., 2020; Laurent & von Brecht, 2017; Bertschinger & Natschläger, 2004; Bertschinger et al., 2004), although previous work did not leverage Lyapunov exponents..
+For S4 (Gu et al., 2022) SSMs, Goel et al.
+(2022) used Hurwitz matrices to characterize the numerical stability of linear time-invariant (LTI) S4 models.
+Similarly,  Orvieto et al.
+(2023) used the LTI property to derive the eigenvalue decomposition of S4-like models, providing stability conditions by bounding the corresponding eigenvalues (further discussed in Appendix E).
+However, such analysis is not applicable to time-varying models, such as Mamba, nor does it characterize the effects of sensitive dependence on initial conditions (e.g., divergence of two $\varepsilon$ close inputs).
+To the best of our knowledge, no previous works have used Lyapunov exponents to explore the effects of mixed-precision on recurrent neural models or Mamba architectures.
+In-context learning.
+Recent work has sought to understand the in-context learning (ICL) capabilities of Mamba LLMs when trained from scratch for specific tasks (Park et al., 2024; Lee et al., 2024).
+Another line of recent work has sought to understand how hybrid Mamba-Transformer models may be directly distilled from Transformer models (Wang et al., 2024b).
+However, to the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics or empirically explored such effects downstream impact on fine-tuning.
+As previously noted, recent works (Park et al., 2024; Lee et al., 2024) have studied Mamba’s ability to perform ICL by training Mamba models for specific tasks.
+Such tasks include logistic regression, decision trees, and learning other simple function classes, following the work of Garg et al.
+(2022).
+We emphasize that, in this set up, relatively small Mamba models–33 million and 90 million parameters for Lee et al.
+(2024) and Park et al.
+(2024), respectively–are trained from scratch for every evaluated task.
+Indeed, Park et al.
+(2024) notes that subsequent work is necessary to understand Mamba’s ICL capabilities for language modeling using standard natural language benchmarks, as well as for larger model sizes.
+Thus, our study of both the pretrained and instruction tuned ICL capabilities of Mamba/Mamba-2 LLMs for natural language tasks are complimentary to previous works.
+Failure Cases and Limitations
+While we explored the use of LoRA for Mamba models, many other PEFT adapters exist (Liu et al., 2022; Li & Liang, 2021; Houlsby et al., 2019; Lester et al., 2021).
+Furthermore, while mixed-precision using FP16 and BF16 were explored, lower-precision methods exist (Dettmers et al., 2024).
+Both are interesting directions for future work.
+Finally, our timing and memory usage experiments using Alpaca did not consider the largest two Mamba models (1.4B and 2.8B) due to their exceeding A10G memory capacity for FP32 full fine-tuning.
+Broader Impact Statement
+The Mamba models considered are all LLMs, and thus have the same potential positive and negative societal impacts as other LLMs (e.g., hallucinations).
+Furthermore, fine-tuning is known to possibly erode existing LLM guardrails, and thus the described methods may be adapted for this fine-tuning use case (as is the case for all PEFT and MPFT methods).
+However, MPFT/PEFT is shown to improve the quality of Mamba models for downstream applications, which may be adapted for all positive LLM applications in society (e.g., personal assistants, task automation, code completion, etc.).
+Finally, the described methods decrease the computational constraints required to train and inference Mamba SSMs, which has implications for green ML (e.g., decreased CO2 emissions, positive climate change impact, etc.).
+Acknowledgments
+We thank Leidos for funding this research through the Office of Technology.
+This manuscript has been approved for public release 24-LEIDOS-0515-27826.
+We thank (NeurIPS) Reviewer Uruo, (TMLR) Reviewer z2sE, and (TMLR) Reviewer SpRc for helpful comments and feedback, which spurred the authors to derive and add several additional theoretical results.
+This work is dedicated to the memory of Bella Halloran; thank you for supporting this one final paper, you are missed.
+Appendix A Mamba stable dynamics proof
 Recall the state-space parameters and equations for the MambaBlock; $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}=[T]$.
@@
 $\bm{x}_{t}$ $=\bar{\mathbf{A}}_{t}\bm{x}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{u}_{t}$ (4)
-$\mathbf{y}_{t}$ $=\bar{\mathbf{C}}_{t}\bm{x}_{t},$
+$\mathbf{y}_{t}$ $=\mathbf{C}_{t}\bm{x}_{t},$
 where $\bar{\bm{\Delta}}_{t}=\texttt{softplus}(\texttt{Linear}(\bm{\Delta}_{t}))\in\mathbb{R}_{\geq 0}{}^{d\times d}$, $\bar{\mathbf{A}}_{t}=\exp{(\bar{\bm{\Delta}}_{t}\mathbf{A})}$, $\bar{\mathbf{B}}_{t}=\mathbf{A}^{-1}(\bar{\mathbf{A}}-\mathbf{I})\mathbf{B}_{t}$, and $\mathbb{R}_{\geq 0}{}$ is the set of non-negative real numbers.
@@
 For input $(\bm{x}_{0},\mathbf{u}_{1})$ to a MambaBlock, small changes $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time.
-That is, $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
+That is, $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
 Proof.
@@
 Recall that for the maximum deviation, we have:
-$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\lambda_{\mathtt{max}}{})}).$
+$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\lambda_{\mathtt{max}}{})}).$
 where the maximal Lyapunov exponent $\lambda_{\mathtt{max}}{}$ is defined as:
@@
 ∎
-Theorem 2.
+Theorem.
 Let $(\bm{x}_{t-1},\mathbf{u}_{t})$ be the latent state and input at an arbitrary time $t\in[1,T]$ within a MambaBlock.
-Then small changes $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially decreasing over discrete-time, i.e., $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
+Then small changes $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially decreasing over discrete-time, i.e., $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
 Proof.
@@
 ∎
-Appendix C Mamba stable outputs proof
-Theorem 3.
+Appendix B Mamba stable outputs proof
+Theorem.
 Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time.
@@
 Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time.
-That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
+That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
 We thus have
@@
 where proportionality follows due to the diagonality of $\mathbf{C}_{N}$ and the vector-absolute value. Thus,
-$\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$ $\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$
-Appendix D Experimental Details
-All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16], version 0.4.2.
-Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models.
-All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0.
-All Mamba-2 models were run using mamba-ssm v2.2.2 using Huggingface checkpoints, e.g., state-spaces/mamba-130m for the smallest model.
-For MPFT, Flash Attention 2.0 [9] via flash_attn 2.5.7 was used for Pythia models.
-For FP16 and BF16 inference results, Flash Attention 2.0 was used for both Pythia and OLMo models.
-For OLMo results, the 336B-token checkpoint was used by specifying revision=step80000-tokens336B.
-All Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps.
-Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively.
-For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 2).
-For SLL LoRA, targeted Mamba layers were {x_proj, embeddings, in_proj, out_proj}; x_proj is the large MambaBlock memory buffer which, when targeted by LoRA, regularizes the majority of SSM parameters during fine-tuning through weight tying (Theorem 3).
-Pythia targeted SLL LoRA layers were {dense, embed_in, query_key_value, dense_h_to_4h,dense_4h_to_h}, chosen to balance performance across model sizes.
-All experiments were run using a single-GPU Nvidia A10G (24 GB total memory).
-For Pythia, Mamba, and Mamba-2 ALL LoRA experiments in Figure 3, all models followed the same training and PEFT recipes, save for Mamba-2 2.7B which required a LoRA $r$ dimension of 64 to fit in A10G memory.
+$\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$ $\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$
+Appendix C Mamba stable AMP and/or PEFT dynamics proof
+For an arbitrary MambaBlock, let $\bm{x}_{0},\dots,\bm{x}_{T}$ and $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ be the trajectory of latent and input states under full-precision fine-tuning.
+During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16).
+During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating.
+MFPT thus introduces changes to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish) and latent states $\bm{x}_{0},\dots,\bm{x}_{T}$ through $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), up/downcasting of parameter updates, and the changes to the aforementioned changes to the input states.
+Potentially in concert with MPFT, PEFT via LoRA induces changes by learning low-rank matrices $\hat{\mathbf{A}}_{t},\hat{\mathbf{B}}_{t},\hat{\mathbf{C}}_{t}$
+$\hat{\bm{x}}_{t}$ $=(\bar{\mathbf{A}}_{t}+\hat{\mathbf{A}}_{t})\bm{x}_{t-1}+(\bar{\mathbf{B}}_{t}+\hat{\mathbf{B}}_{t})\mathbf{u}_{t}$
+$\hat{\mathbf{y}}_{t}$ $=(\mathbf{C}_{t}+\hat{\mathbf{C}}_{t})\bm{x}_{t}$
+Let $\bm{x}_{0}^{\prime},\dots,\bm{x}_{T}^{\prime}$ and $\mathbf{u}_{1}^{\prime},\dots,\mathbf{u}_{T}^{\prime}$ be the trajectory of latent and input states under MPFT and/or PEFT.
+We thus have the following
+Theorem.
+Let $\epsilon_{t}^{u}=\max|\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime}|$, i.e., the maximum, positive-scalar difference between the $t$th MPFT and full fine-tuning input state.
+Similarly, let $\epsilon_{t}^{x}=\max|\bm{x}_{t}-\bm{x}_{t}^{\prime}|$ and $\epsilon^{*}=\max\bigcup_{t=0}^{T}\{\epsilon_{t}^{u},\epsilon_{t}^{x}\}$ (where we define $\mathbf{u}_{0}=0$).
+We thus have
+$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\zeta)}),$
+for some scalar $\zeta\leq 0$. Thus, differences introduced by AMP and/or PEFT for Mamba models do not exponentially compound over discrete-time.
+Proof.
+Consider the $\bm{\epsilon}$-vector definition of the Lyapunov exponent: given $\lambda_{\mathtt{max}}{}$, $\bm{\epsilon}\in\mathbb{R}^{d}$, and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\bm{\epsilon},\mathbf{u}_{1}+\bm{\epsilon})$, the maximum deviation between these points grows as:
+$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}+\bm{\epsilon},\mathbf{u}_{1}+\bm{\epsilon})|\in\mathcal{O}(\left\lVert\bm{\epsilon}\right\rVert_{2}\exp{(T\lambda_{\mathtt{max}}{})}).$
+where $\left\lVert\right\rVert_{2}$ denotes the L2-norm for vectors.
+Let $f(\cdot)$ be a function which maps a vector to the set of its elements, e.g., for $\bm{z}\in\mathbb{R}^{d},f(\bm{x})=\{x_{1},\dots,x_{d}\}$.
+Thus, note that for $a^{*}=\max\{|a|:\forall a\in f(\bm{\epsilon})\}$, $\left\lVert\bm{\epsilon}\right\rVert_{2}\leq a^{*}\sqrt{d}$ and $\mathcal{O}(\left\lVert\bm{\epsilon}\right\rVert_{2}\exp{(T\lambda_{\mathtt{max}}{})})\in\mathcal{O}(a^{*}\exp{(T\lambda_{\mathtt{max}}{})})$.
+Let $\mathcal{E}=\bigcup_{t=0}^{T}\{\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime},\bm{x}_{t}-\bm{x}_{t}^{\prime}\}$ (where we define $\mathbf{u}_{0}=0$), and $\epsilon^{*}=\max\{|a|:\forall a\in\bm{\epsilon},\forall\bm{\epsilon}\in\mathcal{E}\}$.
+$\forall\;\epsilon\in\mathcal{E}$, we thus have
+$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}+\epsilon,\mathbf{u}_{1}+\epsilon)|\in\mathcal{O}(\left\lVert\bm{\epsilon}\right\rVert_{2}\exp{(T\lambda_{\mathtt{max}}{})})\subseteq\mathcal{O}(\epsilon^{*}\exp{(T\lambda_{\mathtt{max}}{})}),$
+Furthermore, $\{\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime},\bm{x}_{t}-\bm{x}_{t}^{\prime}\}\subseteq\mathcal{E}$, and thus
+$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}-\bm{x}_{0},\mathbf{u}_{1}-\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\lambda_{\mathtt{max}}{})})$
+Let $\tau(t)$ be a function that maps time values such that $\tau(t)\in[1,T-t]$ and $\tau(t)=1,\tau(t+1)=2,\dots,\tau(t+T)=T-t$.
+Then $\mathbf{B}_{\tau(t)},\mathbf{C}_{\tau(t)},\bm{\Delta}_{\tau(t)}$ define a new MambaBlock with inputs $\mathbf{u}_{\tau(t)},\dots,\mathbf{u}_{\tau(t+T)}$ and subsequent recurrent states $\bm{x}_{\tau(t)-1},\bm{x}_{\tau(t)},\dots,\bm{x}_{\tau(t+T)}$.
+$\forall t\in\{1,\dots,T\}$, $\{\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime},\bm{x}_{t}-\bm{x}_{t}^{\prime}\}\subseteq\mathcal{E}$, so that
+$\max|F_{\theta}^{T-t}(\bm{x}_{\tau(t)}-1,\mathbf{u}_{\tau(t)})-F_{\theta}^{T-t}(\bm{x}_{\tau(t)-1}^{\prime},\mathbf{u}_{\tau(t)})|\in\mathcal{O}(\epsilon^{*}\exp{((T-t)\lambda_{\mathtt{max}}{})}).$
+Thus, we have
+$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\lambda_{\mathtt{max}}{})}),$
+where $\lambda_{\mathtt{max}}{}\leq 0$ by Lemma 1.
+Appendix D MambaBlock Maximum Deviation results
+D.1 $\varepsilon$ varying initial conditions
+Figure 5: Perturbing the initial states of random MambaBlocks produces output state deviations which exponentially decrease over discrete time.
+$\mu$ denotes the maximum absolute deviation at discrete time $t$ averaged over all trials, while $\sigma$ denotes the respective standard deviation at each discrete time.
+Figure 6: Perturbing the initial states of random MambaBlocks produces output state deviations which exponentially decrease over discrete time.
+$\mu$ denotes the maximum absolute deviation at discrete time $t$ averaged over all trials, while $\sigma$ denotes the respective standard deviation at each discrete time.
+Appendix E Previous SSM stability results
+Previous SSM stability results (Goel et al., 2022; Orvieto et al., 2023) consider the following linear time-invariant (LTI) equations used in S4 models,
+$\bm{x}_{t}$ $=\bar{\mathbf{A}}\bm{x}_{t-1}+\bar{\mathbf{B}}\mathbf{u}_{t}$ (5)
+$\mathbf{y}_{t}$ $=\mathbf{C}\bm{x}_{t}.,$ (6)
+which can be unrolled to produce
+$\bm{x}_{k}$ $=\sum_{j=0}^{k}\bar{\mathbf{A}}^{j}\bar{\mathbf{B}}\mathbf{u}_{j}.$ (7)
+By eigendecomposition,
+$\bar{\mathbf{A}}^{j}$ $=(\mathbf{P}\mathbf{\Lambda}\mathbf{P}^{-1})^{j}=\mathbf{P}\mathbf{\Lambda}^{j}\mathbf{P}^{-1}.$
+In  (Orvieto et al., 2023), the derived sufficient condition for stability is thus $|\alpha_{i}|\leq 1,$ for $\mathbf{\Lambda}=\texttt{diag}([\alpha_{0},\alpha_{1},\dots,\alpha_{d}])$.
+Relatedly, in  (Goel et al., 2022), stability is enforced by constraining $\mathbf{\Lambda}$ to be Hurwitz, in which case the real part of each diagonal entry $\alpha_{i}$ is negative, i.e., $\operatorname{Re}(\alpha_{i})<0,$ for $\mathbf{\Lambda}=\texttt{diag}([\alpha_{0},\alpha_{1},\dots,\alpha_{d}])$.
+Both results critically rely on the LTI property of Equations 5 and  6 to derive Equation 7.
+However, Mamba’s state-space equations are linear time-varying (LTV), and thus Equation 7 does not hold.
+Thus, previous SSM stability results do not directly apply to Mamba models.
+Appendix F Experimental Details
 Table 1: Learning rate and LoRA dimension $r$ values
@@
 2.8B 2.7B 2.8B 5.0e-7 128 64
+All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI (Gao et al., 2023), version 0.4.2.
+Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models.
+All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0.
+All Mamba-2 models were run using mamba-ssm v2.2.2 using Huggingface checkpoints, e.g., state-spaces/mamba-130m for the smallest model.
+For MPFT, Flash Attention 2.0 (Dao et al., 2022) via flash_attn 2.5.7 was used for Pythia models.
+For FP16 and BF16 inference results, Flash Attention 2.0 was used for both Pythia and OLMo models.
+For OLMo results, the 336B-token checkpoint was used by specifying revision=step80000-tokens336B.
+All Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from (Tunstall et al., 2023)): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps.
+Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively.
+For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 1).
+For SLL LoRA, targeted Mamba layers were {x_proj, embeddings, in_proj, out_proj}; x_proj is a large MambaBlock memory buffer which, when targeted by LoRA, adapts parameters $\bm{\Delta}_{t},\mathbf{B}_{t}$, and $\mathbf{C}_{t}$.
+Pythia targeted SLL LoRA layers were {dense, embed_in, query_key_value, dense_h_to_4h,dense_4h_to_h}, chosen to balance performance across model sizes.
+All experiments were run using a single-GPU Nvidia A10G (24 GB total memory).
+For Pythia, Mamba, and Mamba-2 ALL LoRA experiments in Figure 4, all models followed the same training and PEFT recipes, save for Mamba-2 2.7B which required a LoRA $r$ dimension of 64 to fit in A10G memory.
 The Alpaca dataset is freely available for download at https://huggingface.co/datasets/tatsu-lab/alpaca under open-source license CC-by-NC 4.0.
 The OpenHermes dataset is freely available for download at https://huggingface.co/datasets/teknium/OpenHermes-2.5 under open-source license MIT, Apache 2.0, CC.
-Appendix E Mamba-2 mixed-precision inference pretrained LLM performance tables
-Table 3: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference.
-Divergence is averaged over {0, 1, 3, 5}-shot performance.
-Pretrained checkpoints are used for Mamba (M), Mamba-2 (M2), Pythia (P), OLMo [19] (O), and Phi-1.5 [33] (Phi) models.
-Model M M2 P M M2 P M M2 P O M M2 P Phi M M2 P
-Size 130m 130m 160m 370m 370m 410m   790m 780m 1b 1.4b 1.3b 1.4b 1.5b 2.8b 2.7b 2.8b
-FP16 $\mu$ 0.03 0.07 0.35 0.05 0.05 0.06 0.21 0.12 0.05 0.04 0.04 0.04 0.07 0.03 0.15 0.26 0.12
-BF16 $\mu$ 0.05 0.67 1.45 0.20 0.52 0.20 0.66 0.29 0.16 0.13 0.31 0.40 0.13 1.05 1.17 1.02 0.11
-Appendix F Adapted PEFT-layer effects on Mamba-2 zero-shot performance
-As described in Section 3.2, both Mamba and Mamba-2 store the majority of time-varying SSM parameters in a large memory buffer (instantiated as a PyTorch linear layer’s weight matrix), denoted as $\mathbf{W}$.
-For PEFT via LoRA, we ablate the impact of targeting $\mathbf{W}$, as well as the impact of Theorem 3, where time-varying variables accessed through $\mathbf{W}$ were shown to be regularized through weight tying.
-Firstly, we instruction tune Mamba-2 models using LoRA by both adapting $\mathbf{W}$ and adapting all linear layers other than $\mathbf{W}$.
-Mamba-2 MambaBlocks contain two linear layers, thus the former targets only the $\mathbf{W}$ linear layer whereas the latter targets the other linear layer in each block.
-As shown in Table 4, PEFT with only $\mathbf{W}$ targeted within the MambaBlock near uniformly results in better performance than only targeting other layers, with the former outperforming the latter on 32 out of 35 natural language tasks.
-This makes intuitive sense since $\mathbf{W}$ represents the majority of Mamba’s time-varying parameters (as previously mentioned).
-Thus, while previous works [21, 11] have displayed the importance of Mamba’s time-varying parameters for pretraining performance, this verifies the importance of Mamba’s time-varying parameters for instruction tuning performance.
-Table 4: Zero-shot performance for instruction tuned Mamba-2 models where: denotes the large memory buffer $\mathbf{W}$ containing the majority of temporal variables (described in Section 3.2) is targeted for LoRA adaptation, and denotes $\mathbf{W}$ is not adapted.
-The top-performance for each task per model is highlighted in bold .
-Model $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
-ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
-Mamba-2 130M 15.37 45.16 35.41 65.02 47.94 24.91 52.17
-16.95 43.30 35.24 64.85 47.47 24.06 52.17
-Mamba-2 370M 8.03 54.16 46.87 69.53 53.49 27.73 57.14
-8.50 53.74 46.59 70.67 54.80 26.79 55.64
-Mamba-2 780M 5.79 61.61 55.14 71.98 61.11 29.18 60.09
-5.86 61.63 54.94 72.03 60.98 28.41 60.22
-Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
-5.05 65.44 59.84 73.39 64.10 33.11 61.09
-Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
-4.10 69.61 66.60 76.39 69.53 36.26 63.93
-Next, we present a counterintuitive experiment to demonstrate the impact of Theorem 3.
-As before, we instruction tune Mamba-2 models using LoRA, focusing on the two largest model sizes.
-This time, however, we target all linear layers within Mamba-2’s MambaBlock and contrast this with targeting only $\mathbf{W}$ within the MambaBlock.
-Displayed in Table 5, we can see that adapting more SSM parameters does not necessarily lead to improved performance across the board.
-Rather, adapting only $\mathbf{W}$ for Mamba-2 outperforms adapting all linear layers on eight of the 14 natural language tasks.
-The results in Table 5 thus present two cases: (a) regularization: only adapting parameters which lead to weight tying (by Theorem 3) and (b) increased learning capacity: adapting more parameters, at the cost of learning unregularized parameters.
-Through regularization via weight tying [42], the former leads to good generalization.
-In contrast, while adapting more parameters, (b) includes unregularized variables, leading to positive improvements on only a minority of tasks when compared to the fine-tuning of fewer, but regularized, parameters.
-We note that adapting all linear layers is often performed when fine-tuning attention-based LLMs using LoRA [12].
-Thus, a similar regularization-vs-capacity tradeoff for Transformers LLMs may present itself by carefully studying LoRA’s affects when targeting specific linear layers in attention-based architectures.
-Table 5: Zero-shot performance for instruction tuned Mamba-2 models.
-denotes only $\mathbf{W}$ was targeted for LoRA adaptation within the MambaBlock, and thus denotes both linear layers within each MambaBlock were adapted.
-The top-performance for each task per model is highlighted in bold.
-Model Only $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
-ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
-Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
-4.94 62.95 60.06 73.78 64.02 34.81 61.33
-Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
-3.77 68.39 67.51 75.95 69.53 38.31 65.35
+Table 2: Pretrained model performance.
+Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI (Gao et al., 2023).
+LAMBADA zero-shot is more effective for the model sizes considered (further discussed in (Xie et al., 2021; Brown et al., 2020)) and thus excluded from few-shot performance averages.
+Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
+Model NN-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
+ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
+Mamba 130M 0 16.0 44.3 35.2 64.7 48.0 24.3 52.6
+1 19.3 38.2 35.1 64.6 47.0 23.5 50.8 -1.9
+3 23.1 35.2 35.0 65.1 49.1 24.0 51.0 -0.4
+5 24.4 36.2 34.9 64.9 49.1 23.7 50.0 -1.2
+Mamba-2 130M 0 16.8 43.9 35.3 64.9 47.4 24.2 52.6
+1 20.6 37.9 34.9 64.1 46.9 23.1 51.3 -2.0
+3 24.3 35.1 34.9 64.4 49.0 24.7 52.9 0.9
+5 26.5 34.9 34.6 64.4 48.6 24.8 51.7 0.2
+Pythia 160M 0 38.2 32.7 30.2 61.8 43.4 23.8 51.0
+1 47.2 28.2 30.6 62.2 43.4 23.7 49.3 -0.4
+3 63.7 24.7 30.5 61.9 44.8 22.9 51.3 0.1
+5 66.3 25.3 30.4 62.6 43.4 23.1 50.8 -0.2
+Mamba 370M 0 8.1 55.6 46.5 69.5 54.9 27.8 55.3
+1 9.7 49.8 45.9 69.3 57.4 26.5 54.7 -0.5
+3 10.9 48.4 46.2 69.5 58.8 28.4 53.8 1.2
+5 11.4 48.6 46.2 69.4 58.3 28.0 55.9 1.5
+Mamba-2 370M 0 8.0 55.9 46.9 70.5 54.8 26.7 55.4
+1 9.8 50.3 46.4 70.5 56.5 26.8 54.2 0.0
+3 11.3 48.5 46.6 70.2 59.0 26.9 54.3 1.0
+5 12.5 46.6 46.7 70.3 58.5 28.2 53.3 1.5
+Pythia 410M 0 10.8 51.5 40.6 66.9 52.0 24.1 53.4
+1 12.3 47.1 40.5 68.0 53.8 25.6 52.4 1.8
+3 14.4 43.2 40.9 67.9 55.1 26.9 54.0 4.2
+5 14.6 44.1 40.8 68.1 54.6 26.6 53.4 3.5
+Mamba 790M 0 6.0 61.4 55.1 72.3 61.2 29.5 55.9
+1 7.1 55.9 54.5 72.4 63.0 30.2 56.9 1.3
+3 8.1 54.5 54.2 72.3 63.5 31.4 57.1 2.2
+5 8.8 52.9 54.6 72.6 64.4 31.9 57.2 3.1
+Mamba-2 780M 0 5.9 61.7 54.9 72.0 61.0 28.5 60.2
+1 7.1 55.5 54.7 72.4 62.3 32.1 57.1 1.9
+3 8.6 53.3 54.7 72.5 62.8 32.3 57.8 2.5
+5 9.9 51.4 55.2 72.1 62.8 32.2 56.8 2.2
+Pythia 1B 0 7.9 56.3 47.2 70.7 57.0 27.0 53.4
+1 8.0 51.8 47.3 70.7 57.1 28.2 53.4 1.0
+3 10.5 48.2 47.5 71.2 59.2 28.0 54.3 2.2
+5 10.9 48.4 47.3 71.4 58.7 28.4 53.1 1.9
+Mamba 1.4B 0 5.0 64.9 59.2 74.1 65.5 32.9 61.3
+1 5.8 60.6 58.2 74.7 64.5 33.0 60.9 -0.5
+3 6.6 58.9 58.9 73.6 66.1 34.5 60.9 0.7
+5 7.0 58.3 59.0 74.1 66.4 35.5 60.4 1.5
+Mamba-2 1.3B 0 5.0 65.6 60.0 73.2 64.2 33.1 61.1
+1 6.0 60.1 59.4 73.1 65.6 35.3 59.4 1.0
+3 6.7 58.6 60.1 73.4 66.5 35.4 61.9 2.5
+5 7.0 58.6 60.2 73.7 66.5 35.9 61.4 2.7
+Pythia 1.4B 0 6.1 61.7 52.1 70.9 60.5 28.5 57.4
+1 7.0 56.3 52.1 71.4 62.0 29.5 57.5 1.4
+3 7.9 54.4 52.6 70.9 63.9 31.1 56.8 2.9
+5 8.0 54.4 52.8 71.0 63.2 31.3 57.8 3.3
+Mamba 2.8B 0 4.2 69.1 66.1 75.2 69.6 36.4 63.3
+1 5.0 63.7 65.6 75.6 69.9 37.1 63.9 0.6
+3 5.5 62.8 65.5 75.3 70.8 38.1 65.1 1.7
+5 5.7 62.5 66.1 76.1 70.9 38.1 64.6 2.0
+Mamba-2 2.7B 0 4.1 69.6 66.6 76.4 69.5 36.3 63.9
+1 4.8 65.1 65.9 75.1 70.0 38.6 65.1 1.3
+3 5.3 63.9 66.8 75.2 71.9 41.0 64.1 3.1
+5 5.7 62.3 67.1 75.3 70.7 41.2 65.9 3.6
+Pythia 2.8B 0 5.0 64.7 59.3 73.9 64.2 32.9 59.8
+1 5.7 60.9 59.4 73.8 66.8 34.8 59.0 1.7
+3 6.2 59.1 59.9 74.7 67.4 34.9 60.8 2.9
+5 6.5 59.1 60.2 74.5 67.1 35.0 61.3 3.1
+Table 3: Instruction tuned model performance.
+Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI (Gao et al., 2023).
+LAMBADA zero-shot is more effective for the model sizes considered (further discussed in (Xie et al., 2021; Brown et al., 2020)) and thus excluded from few-shot performance averages.
+Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
+Model NN-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
+ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
+Mamba 130M 0 12.9 46.5 35.1 64.2 48.7 25.5 51.7
+1 17.8 38.1 35.0 64.2 48.6 24.9 52.2 -0.4
+3 22.3 35.3 34.8 64.2 50.2 24.5 50.6 -0.8
+5 23.6 35.9 34.7 64.7 49.8 24.6 50.2 -0.9
+Mamba-2 130M 0 15.2 44.5 35.1 64.5 47.2 24.7 52.2
+1 21.9 36.1 34.5 64.3 46.8 24.0 50.8 -1.7
+3 26.9 33.3 34.7 65.1 48.5 25.2 51.5 0.6
+5 29.0 33.8 34.5 64.8 48.7 25.1 51.3 0.4
+Pythia 160M 0 30.2 36.1 30.0 62.2 44.7 23.6 50.3
+1 44.5 29.1 30.4 62.0 44.0 23.6 50.5 -0.0
+3 66.7 25.5 30.3 62.8 45.2 22.8 49.8 -0.3
+5 70.4 25.3 30.5 62.9 44.1 23.4 50.8 0.3
+Mamba 370M 0 7.2 56.0 46.3 69.2 55.3 27.7 56.0
+1 9.3 49.9 45.7 68.7 57.1 28.3 55.4 0.5
+3 10.4 49.4 45.7 68.9 58.7 29.7 54.1 1.6
+5 11.0 48.3 45.7 70.1 59.3 29.1 54.5 1.9
+Mamba-2 370M 0 7.6 54.7 46.8 69.3 52.2 27.0 56.0
+1 9.9 48.3 46.0 69.6 55.7 28.8 55.2 2.1
+3 11.8 46.3 46.3 70.1 59.0 29.1 54.5 3.6
+5 12.6 45.5 46.3 70.8 59.6 29.5 53.0 3.8
+Pythia 410M 0 13.3 46.4 40.9 67.4 52.7 25.4 53.4
+1 17.2 40.4 40.5 68.4 53.6 25.7 53.0 0.5
+3 21.1 37.4 40.9 67.7 55.7 27.1 52.6 2.3
+5 21.5 38.2 40.7 67.8 55.7 27.3 53.8 2.8
+Mamba 790M 0 5.2 62.8 55.6 72.8 62.4 30.6 56.2
+1 6.3 56.6 54.9 72.7 64.6 31.7 56.3 1.2
+3 7.0 55.6 54.7 72.4 65.3 33.2 57.5 2.7
+5 7.5 54.6 54.9 72.9 65.6 33.8 57.2 3.2
+Mamba-2 780M 0 4.9 63.4 55.8 71.7 61.1 30.6 59.2
+1 6.6 55.2 54.4 72.7 64.2 34.0 57.6 2.5
+3 7.8 52.7 54.9 73.5 65.0 34.6 57.8 3.6
+5 8.6 52.8 54.8 73.4 64.6 34.0 58.0 3.1
+Pythia 1B 0 7.7 56.6 47.3 70.8 57.1 26.7 53.4
+1 8.8 52.0 47.4 70.7 57.5 28.8 53.6 1.8
+3 10.2 48.7 47.5 71.4 59.0 28.5 54.4 2.6
+5 10.6 48.8 47.4 71.5 58.9 28.4 53.0 2.0
+Mamba 1.4B 0 4.6 64.8 59.3 74.3 65.2 35.1 62.3
+1 5.4 60.3 58.2 74.3 66.7 35.7 62.8 0.6
+3 6.1 59.3 58.4 74.1 67.4 36.6 61.8 1.0
+5 6.3 58.8 58.8 74.5 68.3 37.0 59.9 1.1
+Mamba-2 1.3B 0 4.9 63.0 60.1 73.8 64.0 34.8 61.3
+1 6.1 58.2 59.2 74.2 67.0 35.0 60.1 0.5
+3 7.0 56.6 59.4 73.7 67.8 36.6 59.9 1.5
+5 7.2 56.5 59.9 73.5 68.5 36.7 60.7 2.2
+Pythia 1.4B 0 5.2 63.6 52.9 71.1 61.2 30.3 58.2
+1 6.2 57.4 52.7 71.7 62.2 30.6 56.9 0.2
+3 7.0 56.1 53.1 71.1 64.5 32.8 56.8 2.3
+5 7.1 55.5 53.3 71.2 63.8 33.5 57.5 2.9
+Mamba 2.8B 0 4.0 67.7 66.4 75.6 68.4 36.6 64.2
+1 4.8 63.3 65.9 76.2 70.9 39.4 64.6 2.4
+3 5.3 62.1 65.7 75.8 71.3 39.1 65.4 2.4
+5 5.4 61.9 66.2 77.2 71.4 40.4 66.1 3.9
+Mamba-2 2.7B 0 3.8 68.4 67.5 76.0 69.5 38.3 65.3
+1 4.5 63.8 66.7 76.0 71.8 41.5 67.1 2.6
+3 5.0 62.3 67.3 76.2 73.3 44.4 66.0 4.5
+5 5.3 61.8 67.4 76.4 72.4 44.5 65.0 4.1
+Pythia 2.8B 0 5.0 64.7 59.3 74.0 64.7 33.3 59.2
+1 5.6 60.8 59.5 74.0 66.7 34.9 59.3 1.7
+3 6.1 59.2 59.9 75.0 67.5 34.9 60.9 2.9
+5 6.5 59.0 60.4 74.5 67.0 35.1 61.2 3.0
+F.1 Hardware throughput and memory-utilization improvements given PEFT and MPFT
+With stable dynamics and observed divergences smaller than comparable Transformers, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs.
+To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset.
+However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
+For each MPFT and PEFT configuration, the average tokens-per-second (ATPS) is calculated as the total tokens used for fine-tuning divided by total training time, and the maximum memory-per-token (MMPT) is calculated as the maximum GPU memory utilization incurred (over the entire fine-tuning run) divided by the total number of tokens in each mini-batch.
+Results are plotted in Figure 7.
+Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 7.
+Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%.
+Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs.
+Furthermore, while full fine-tuning is no longer possible on a single A10G for Mamba models greater than 790 million parameters, MPFT and PEFT allow training Mamba models up to 2.8 billion parameters on GPUs with as little as 24 GB onboard memory.
+(a) Average tokens-per-second $\uparrow$.
+(b) Maximum memory-per-token $\downarrow$.
+Figure 7: Timing and memory usage calculated Mamba model-sizes and PEFT combinations.
+Each model was trained using the Alpaca dataset dataset for three epochs and maximum sequence length 512.
+For each PEFT combination, the batch size was tuned to maximize GPU occupancy.
+Full fine-tuning exceeds available GPU memory (24 GB) for models greater than 790 million parameters.
+Appendix G ICL as an emergent ability of Mamba SSMs
+We study the emergent behavior (as a function of model size) of Mamba/Mamba-2 SSMs’ ICL abilities on natural language tasks by comparing to a larger number of Transformer-based LLMs of varying sizes.
+We compare to OpenELM (Mehta et al., 2024) (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B (Zhang et al., 2024), and OLMo 1.2B (Groeneveld et al., 2024).
+To limit the emergent effects on both parameter size and pretraining token counts, we did not evaluate models greater than 2.8 billion parameters and chose open-source checkpoints as close as possible to the 300 billion total pretraining tokens used for Mamba, Mamba-2, and Pythia models.
+Thus, pretraining token counts for OpenELM, TinyLlama, and OLMo models were 429 billion, 503 billion, and 336 billion, respectively.
+We note this potentially biases ICL performance in favor of the newly evaluated Transformer-based LLMs, and that direct comparisons between Mamba, Mamba-2, and Pythia are the most fair (as these three classes of models were all pretrained on the same dataset for the same number of total pretraining tokens).
+We repeat the experiments from Figure 4, where we evaluate the pretrained and instruction tuned ICL capabilities of all models.
+To understand the critical role of parameter counts, we group all models into two classes: LLMs containing 450 million parameters or less, and LLMs containing greater than 450 million parameters.
+ICL performance measured by AIPSS is displayed in Figure 8.
+From Figure 8, it is clear that pretrained SSMs and Transformers of parameter counts 270 million and less display slight or detrimental ICL abilities (i.e., few-shot performance is worse than zero-shot).
+For models of greater than 450 million parameters, the majority of SSMs and Transformers display positive ICL abilities, with Mamba 1.4B being an outlier in terms of poor performance.
+With the exception of TinyLlama at 1-shot performance and Mamba-2 2.7B for 3- and 5-shot performance, the majority of other pretrained models cluster together.
+Instruction tuning greatly smooths ICL performance across both parameter classes.
+While instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all parameters of 270 million and greater show positive ICL abilities.
+For instruction tuned models of greater than 450 million parameters, all SSMs and Transformers show positive ICL abilities, with Mamba-2 2.7B greatly outperforming all other models (both SSM and Transformer) in this class.
+Thus, in terms of ICL as a function of SSM model size, while no clear trend presents itself for pretrained models, ICL emerges for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater.
+(a) $450\text{M}<\text{Parameter total}$, Pretrained
+(b) $450\text{M}<\text{Parameter total}$, ALL LoRA
+(c) $\text{Parameter total }\leq 450\text{M}$, Pretrained
+(d) $\text{Parameter total }\leq 450\text{M}$, ALL LoRA
+Figure 8: Instruction tuning improves Mamba-2 ICL performance past Transformer LLMs.
+ALL LoRA models were instruction fine-tuned on the OpenHermes dataset for one epoch.
+Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks: HellaSwag, PIQA, Arc-E, Arc-C, and WinoGrande.
+Mamba* and Mamba’ refer to recent Mamba-1 PEFT methods, i.e., $\text{LoRA}_{p}(X)$ and Additional-Scan, respectively, from Yoshimura et al.
+(2025).
+Appendix H Expanded divergence results: Alpaca and LIMA fine-tuning, MMLU and Winogrande benchmarks, Mean and Standard Deviation divergences
+We extend the non-divergent Mamba fine-tuning results from Section 4. Recall that the following MPFT and PEFT configurations are considered to fine-tune each considered LLM:
+  1. 1.
+    Full fine-tuning in FP32
+  2. 2.
+    Full fine-tuning in FP16
+  3. 3.
+    Full fine-tuning in BF16
+  4. 4.
+    ALL LoRAfine-tuning in FP32
+  5. 5.
+    ALL LoRAfine-tuning in FP16
+  6. 6.
+    ALL LoRAfine-tuning in BF16
+  7. 7.
+    SLL LoRAfine-tuning in FP32
+  8. 8.
+    SLL LoRAfine-tuning in FP16
+  9. 9.
+    SLL LoRAfine-tuning in BF16
+In addition to the Alpaca dataset (Taori et al., 2023), we also fine-tune all models using the LIMA dataset (Zhou et al., 2024).
+Models are trained using LIMA for 5 epochs, while all other settings follow the fine-tuning recipe used for Alpaca (described in Appendix F).
+For natural language benchmarks, in addition to MMLU, we evaluate each fine-tuned model using Winogrande (Sakaguchi et al., 2021).
+Recall that, for each benchmark, divergence between a mixed-precision fine-tuned model is measured between its full-precision counterpart and averaged over {0, 1, 3, 5}-shot performance.
+In addition to the average divergence, we also include the standard deviation of divergence.
+Thus, in total, 144 LLMs were fine-tuned, 576 MMLU evaluations were conducted, and 576 Winogrande evaluations were conducted.
+Figure 9 displays results for Alpaca and MMLU, Figure 10 displays results for Alpaca and Winogrande, Figure 11 displays results for LIMA and MMLU, Figure 12 displays results for LIMA and Winogrande.
+Summary statistics for all experiments are presented in Table 4.
+While OpenELM exhibits large deviation spikes for both Alpaca benchmark evaluations–and Pythia exhibits large deviation spikes for all four evaluations–Mamba does not exhibit a single large deviation spike on any benchmark for all considered model sizes and MPFT/PEFT configurations (i.e., 18 total configurations excluding the full-precision baselines).
+Furthermore, Mamba models are significantly more stable for MPFT/PEFT compared to Transformer-based LLMs.
+E.g., for MMLU evaluations, Alpaca fine-tuning with Mamba models is an average 2.6 times smaller in mean divergence than both Pythia and OpenELM models, while LIMA fine-tuning with Mamba models is an average 7 and 3.25 times smaller in mean divergence than Pythia and OpenELM models, respectively.
+Table 4: Summary of divergence results for Alpaca and LIMA fine-tuning datasets, MMLU and Winogrande benchmarks, and Mamba, OpenELM, and Pythia models.
+For each deviation summary statistic per fine-tuning dataset and benchmark, the lowest deviation is highlighted in bold.
+(Fine-tuning dataset), Benchmark Architecture Large deviation spikes $\downarrow$ Avg mean divergence $\downarrow$ Std mean divergence $\downarrow$
+(Alpaca, MMLU ) Pythia 1 0.37 0.41
+OpenELM 1 0.37 0.32
+Mamba 0 0.14 0.08
+(Alpaca, Winogrande ) Pythia 4 0.72 0.58
+OpenELM 3 0.59 0.37
+Mamba 0 0.25 0.09
+(LIMA, MMLU ) Pythia 1 0.28 0.34
+OpenELM 0 0.13 0.15
+Mamba 0 0.04 0.03
+(LIMA, Winogrande ) Pythia 3 0.45 0.45
+OpenELM 0 0.36 0.18
+Mamba 0 0.11 0.12
+(a) Mean full-precision (FP32) divergence in MMLU performance.
+Figure 9: Alpaca fine-tuning, MMLU evaluation.
+Mamba, Pythia, and OpenELM models are fine-tuned over the Alpaca dataset using different combinations of MPFT and PEFT.
+Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
+(a) Mean full-precision (FP32) divergence in Winogrande performance.
+Figure 10: Alpaca fine-tuning, Winogrande evaluation.
+Mamba, Pythia, and OpenELM models are fine-tuned over the Alpaca dataset using different combinations of MPFT and PEFT.
+Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
+(a) Mean full-precision (FP32) divergence in MMLU performance.
+Figure 11: LIMA fine-tuning, MMLU evaluation.
+Mamba, Pythia, and OpenELM models are fine-tuned over the LIMA dataset using different combinations of MPFT and PEFT.
+Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
+(a) Mean full-precision (FP32) divergence in Winogrande performance.
+Figure 12: LIMA fine-tuning, Winogrande evaluation.
+Mamba, Pythia, and OpenELM models are fine-tuned over the LIMA dataset using different combinations of MPFT and PEFT.
+Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.